### v1

# Differentially private projection-depth-based medians

Abstract
We develop $(\epsilon,\delta)$-differentially private projection-depth-based medians using the propose-test-release (PTR) and exponential mechanisms. Under general conditions on the input parameters and the population measure, (e.g. we do not assume any moment bounds), we quantify the probability the test in PTR fails, as well as the cost of privacy via finite sample deviation bounds. We demonstrate our main result on the canonical projection-depth-based median. In the Gaussian setting, we show that the resulting deviation bound matches the known lower bound for private Gaussian mean estimation, up to a polynomial function of the condition number of the covariance matrix. In the Cauchy setting, we show that the “outlier error amplification” effect resulting from the heavy tails outweighs the cost of privacy. This result is then verified via numerical simulations. Additionally, we present results on general PTR mechanisms and a uniform concentration result on the projected spacings of order statistics.
1 Introduction
Privacy has become a central concern in data collection and analysis owing to considerations around personal autonomy, freedom of expression, and fairness and equality. Differential privacy has emerged as a powerful framework for maintaining privacy in data collection and analysis. It has recently been adopted by many high profile institutions, including the United States Census Bureau [1]. This has spurred investigations of many classical statistical problems through a differentially private lens. Among such problems is median estimation.
Early literature on differentially private median estimation focuses on the univariate setting [10, 3, 7, 24]. In the multivariate setting, several authors have studied differentially private medians through the lens of differentially private mean estimation. These works have mainly focused either on the Tukey median [6, 17, 4], or variants of it [15]. In addition, as part of a more general framework for private high-dimensional estimation, Liu et al., 2021b [18] studied one instance of a projection-depth-based median based on the trimmed mean and the trimmed variance. Given that their focus was mean estimation, all of these works enforce a bound on the number of moments possessed by the population measure. However, a classical method for assessing the robustness of a median is to examine its behavior in the case where the population measure possesses no moments. Recently, Ramsay et al., [20] considered this problem. They studied differentially private multivariate medians arising from depth functions, including the Tukey median, without enforcing any moment bounds on the population measure. However, their framework does not cover projection-depth-based medians [26, 27].
Unlike the Tukey median and those studied by Ramsay et al., [20], when the location and scale measures are chosen appropriately, projection-depth-based medians simultaneously possess affine equivariance and achieve a high breakdown point [26, 27]. Affine equivariance is a critical property for location estimates, as it ensures that the units of measurement do not have an effect on the resulting estimate, see, for example, Zuo and Serfling, [28]. In addition, the breakdown point is a staple measure of the robustness of an estimator. It measures the fraction of points which must be corrupted to make an estimate arbitrarily large [14]. For instance, the breakdown point of the Tukey median degrades quickly with the dimension $(1+d)^{-1}$, whereas, in general, the projection-depth-based medians retain a much higher breakdown point, (e.g., $\approx 1/2-d/n$ [26]). One may also note that (when the location and scale measures are chosen appropriately) the projection-depth-based medians enjoy higher relative efficiency than that of the Tukey median [26]. Motivated by these properties, our present work develops and studies differentially private projection-depth-based medians. We consider a general location estimation setting and do not place any moment bounds on the population measure.
Our main tool for privatizing projection-depth-based medians is the propose-test-release (PTR) algorithm, which was introduced by Dwork and Lei, [10] in the context of differentially private estimation of various robust, univariate statistics. Our work takes inspiration from that work, as well as from the works of Brunel and Avella-Medina, [7] and Brown et al., [6]. Brunel and Avella-Medina, [7] used PTR to develop sub-Gaussian median estimators in the univariate setting. Additionally, Brown et al., [6] developed a private Gaussian mean estimator, using a PTR algorithm based on safety and the exponential mechanism, both of which we incorporate into our proposed algorithm. Notably, Brown et al., ’s algorithm was also generalized by Liu et al., 2021b [18] to be an all-purpose solution to various private, statistical problems. Our work builds on that of Liu et al., 2021b [18], in the sense that their mean estimator fits into the class of projection-depth-based estimators.
Our contributions are as follows: We first propose a class of $(\epsilon,\delta)$-differentially private projection-depth-based medians. Next, under general conditions on the population measure, we quantify both the probability that the PTR algorithm fails to release an estimate and the cost of privatizing the projection-depth-based medians (see Theorems 3.1, and 3.2, respectively). To measure the cost of privacy, we provide finite sample bounds on the distance between a private projection-depth-based median and the corresponding population value, which hold with high probability. We then demonstrate our methods for the canonical projection-depth-based median. In particular, we show that Theorems 3.1 and 3.2 apply provided the “projected densities” are bounded below in an interval containing a range of inner quantiles (Condition 4). As a first example, we show that in the Gaussian setting, our deviation bound matches the optimal lower bound for sub-Gaussian mean estimation given by [16, 8], up to a polynomial function of the condition number of the covariance matrix (see Example 3.1). We then show that in the Cauchy setting, the increase in error resulting from contamination or heavy tails, which we call the “outlier error amplification” is larger than the cost of privacy (see Example 3.2). We then confirm this phenomenon with numerical studies (see Section 4). Lastly, we provide some results for analysis of general PTR mechanisms (see Section 5), as well as a uniform concentration result on the projected spacings of order statistics (see Lemma C.2), which may be of general interest.
2 Preliminaries
2.1 Differential privacy
We first review necessary concepts from differential privacy. For a textbook introduction to differential privacy, see Dwork and Roth, [11]. A dataset of size $n\times d$ is a collection of $n$ points in $\mathbb{R}^{d}$, $\mathbb{X}_{n}=\{X_{i}\}_{i=1}^{n}$, possibly with repetitions. Let $\mathbf{D}_{n\times d}$ be the set of all datasets of size $n\times d$. For datasets $\mathbb{X}_{n}$ and $\mathbb{Y}_{n}$, let $\mathbb{X}_{n}\triangle\mathbb{Y}_{n}$ represent the dataset containing all rows in $\mathbb{X}_{n}$ which are not contained in $\mathbb{Y}_{n}$. Define $\mathbf{A}(\mathbb{X}_{n},m)=\left\{\mathbb{Y}_{n}\in\mathbf{D}_{n\times d}\colon|\mathbb{X}_{n}\triangle\mathbb{Y}_{n}|=m\right\}$, the set of datasets of size $n\times d$ which differ from $\mathbb{X}_{n}$ in exactly $m$ rows. Define the empirical measure of $\mathbb{X}_{n}$, ${\hat{\nu}}_{\mathbb{X}_{n}}$, as that which assigns $1/n$ probability to each row in $\mathbb{X}_{n}$. For a dataset with empirical measure ${\hat{\nu}}_{\mathbb{X}_{n}}$ define
~$\widetilde{\mathcal{M}}_{1}({\hat{\nu}}_{\mathbb{X}_{n}},m)=\{\tilde{\nu}\in\mathcal{M}_{1}(\mathbb{R}^{d})\colon\tilde{\nu}={\hat{\nu}}_{\mathbb{Y}_{n}}\text{ for some }\mathbb{Y}_{n}\in\mathbf{A}(\mathbb{X}_{n},m)\},$
where $\mathcal{M}_{1}(\mathbb{R}^{d})$ is the set of measures on $\mathscr{B}(\mathbb{R}^{d})$, and $\mathscr{B}(S)$ denotes the Borel sets of $S$.
Suppose that we observe a dataset ${\mathbb{X}}_{n}$ comprised of $n$ i.i.d. samples from a measure $\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$, and wish to estimate a parameter $\theta\in\Theta$. One way to do this is via a mechanism. Formally, a mechanism, $P_{\hat{\nu}}$, maps an empirical measure $\hat{\nu}$ to $\mathcal{M}_{1}(\Theta)$. The estimator $\tilde{\theta}$ is then taken to be a draw from $P_{\hat{\nu}}$. When the corresponding measure is clear from context, we will suppress dependence on it in notation, denoting $P_{\hat{\nu}}$ as $P$. To produce a differentially private statistic, $P_{\hat{\nu}}$ should satisfy the following privacy guarantee.
Definition 2.1.
A mechanism $P$ is $(\epsilon,\delta)$-differentially private if for any dataset $\mathbb{X}_{n}$, any $\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\nu_{\mathbb{X}_{n}},1)$, and any measurable set $B\in\mathscr{B}(\Theta)$, it holds that
$P_{\hat{\nu}_{\mathbb{X}_{n}}}(B)\leq e^{\epsilon}P_{\tilde{\nu}}(B)+\delta.$ (2.1)
One particular class of differentially private mechanisms are those based on propose-test-release (PTR) algorithms [10]. PTR algorithms generally rely on the notion of local sensitivity. Colloquially, a statistic, $T$, is locally sensitive with respect to a dataset $\mathbb{X}_{n}$ if $\sup_{\mathbb{Y}_{n}\in\mathbf{A}(\mathbb{X}_{n},m)}|T(\mathbb{X}_{n})-T(\mathbb{Y}_{n})|$ is large for some small $m\in\mathbb{N}$. PTR algorithms are characterized by a three-step procedure. First, you propose a statistic. Second, you test if the proposed statistic is locally sensitive. Finally, if the proposed statistic is not locally sensitive, you release a noisy version of the statistic; if not, nothing is released. Even in the event that nothing is released, the PTR mechanism uses up the $(\epsilon,\delta)$ privacy budget. To formalize PTR, take ^$\widehat{\mathcal{M}}_{1}(n,d)$ to be the set of empirical measures constructed from datasets of size $n\times d$, $\{\nu_{\mathbb{X}_{n}}\colon\ \mathbb{X}_{n}\in\mathbf{D}_{n\times d}\}$. Let $Ber(p)$ denote the Bernoulli measure with parameter $0\leq p\leq 1$.
Definition 2.2.
A test mechanism is any mechanism, $Q$, such that for all $\hat{\nu}\in\widehat{\mathcal{M}}_{1}(n,d)$, $Q_{\hat{\nu}}=Ber(\lambda(\hat{\nu}))$ for some $0\leq\lambda(\hat{\nu})\leq 1$.
Every test mechanism is characterized by its associated $\lambda$. We will refer to a test mechanism by its associated $\lambda$ function.
Definition 2.3.
Given a mechanism $Q$ mapping to $\mathcal{M}_{1}(\Theta)$ and a test mechanism $\lambda$, the PTR mechanism based on $(Q,\lambda)$ is given by the map
$P_{\hat{\nu}}\colon(B)\mapsto\lambda(\hat{\nu})Q_{\hat{\nu}}(B\backslash\{\perp\})+\left[1-\lambda(\hat{\nu})\right]\mathbbm{1}\left\{\perp\in B\right\},$
where $B\in\mathscr{B}(\Theta\cup\{\perp\})$ and $\hat{\nu}\in\widehat{\mathcal{M}}_{1}(n,d)$.
The probability of any event under the PTR mechanism will be the probability that the test is passed multiplied by the probability of the event under the original mechanism, plus the probability of the test failing if $\perp$ is an element of the event. We now introduce the notions of safety, safety margin, and the exponential mechanism, which will be used to define the private projection-depth-based median.
Safety, introduced by Brown et al., [6] for Gaussian mean estimation, represents a localized version of the $(\epsilon,\delta)$-differential privacy guarantee. We present a generalized definition.
Definition 2.4.
Given $\epsilon,\delta>0$, the pair $(Q,\hat{\nu})$ is $(\epsilon,\delta)$-safe if, for all $\tilde{\nu}\in\widetilde{\mathcal{M}}(\hat{\nu})$ and $Q$-measurable sets $B$, it holds that
$Q_{\hat{\nu}}(B;\epsilon,\delta)\leq e^{\epsilon}Q_{\tilde{\nu}}(B;\epsilon,\delta)+\delta\qquad\text{and}\qquad Q_{\tilde{\nu}}(B;\epsilon,\delta)\leq e^{\epsilon}Q_{\hat{\nu}}(B;\epsilon,\delta)+\delta.$
The $(\epsilon,\delta)$-safe set of $Q$ is
[…]
We include $\epsilon$ and $\delta$ in the notation for the mechanism $Q$ to emphasize that $Q$ typically depends on $\epsilon$ and $\delta$. Our proposed algorithm’s test mechanism relies on the safety margin.
Definition 2.5.
Given $\epsilon,\delta>0$ and a mechanism $Q$, the safety margin of an empirical measure $\hat{\nu}$ with respect to $Q$, is
$\operatorname{SM}(\hat{\nu},Q;\epsilon,\delta)=\inf\{m\in\mathbb{Z}^{+}\colon\ \exists\tilde{\nu}\in\tilde{\mathcal{M}}(\hat{\nu},m)\ s.t.\ \tilde{\nu}\notin\mathbf{S}(Q,\epsilon,\delta)\}.$
When $\epsilon$ and $\delta$ are clear from the context, we omit them, denoting the safety margin as $\operatorname{SM}(\hat{\nu},Q).$ Observe that if $\operatorname{SM}(\hat{\nu},Q)>0$, then $(Q,\hat{\nu})$ is safe.
Our proposed algorithm applies a test mechanism based on the safety margin, and if the test is passed, releases a draw from the exponential mechanism. For a cost function $\phi(\cdot;\hat{\nu})$, let $A_{\tau,\hat{\nu}}=\{\phi(x,\hat{\nu})\leq\tau\}$. The exponential mechanism $\operatorname{EM}_{\hat{\nu},\phi}(\cdot;\tau,\eta,\epsilon)$ is the measure whose Radon–Nykodym density satisfies
$d\operatorname{EM}_{\hat{\nu},\phi}(\cdot;\tau,\eta,\epsilon)\propto\mathbbm{1}\left\{x\in A_{\tau,\hat{\nu}}\right\}\exp(-\phi(x,{\hat{\nu})\epsilon/4\eta})dx.$ (2.2)
We omit the parameters $(\tau,\eta,\epsilon)$ in the notation when they are clear from the context.
2.2 Outlyingness and projection-depth-based medians
We propose using the projected outlyingness [9, 28, 26] for the cost function of the exponential mechanism. For a measure $\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$ and $u$ from the unit $(d-1)$-sphere, $\mathbb{S}^{d-1}$, let $\nu_{u}$ be the law of $X^{\top}u$ if $X\sim\nu$.
Definition 2.6.
Given a translation and scale equivariant estimator $\mu$ and a translation invariant and scale equivariant estimator $\sigma$, the projected outlyingness at $x\in\mathbb{R}^{d}$ with respect to $\nu$ is
$O(x,\nu)=\sup_{u\in\mathbb{S}^{d-1}}O_{u}(x,\nu_{u})=\sup_{u\in\mathbb{S}^{d-1}}\frac{\left|x^{\top}u-\mu(\nu_{u})\right|}{\sigma(\nu_{u})}.$ (2.3)
[…]
The projection-depth-based medians enjoy several nice properties: local and global robustness, affine equivariance, strong $\sqrt{n}$-consistency and a relatively simple limiting distribution [26, 27]. Notably, to the best of our knowledge, among the competing affine equivariant medians, when the location and scale measures are chosen appropriately, e.g. the median and median absolute deviation, the projection-depth-based medians have the highest breakdown point [26]. For instance, the breakdown point of the Tukey median degrades quickly with the dimension $(1+d)^{-1}$, whereas, in general, the projection-depth-based medians retain a much higher breakdown point, e.g., $\lfloor n/2-d+2\rfloor/n$ [26]. One may also note that the projection-depth-based medians have been shown to have higher asymptotic efficiency, relative to competitors [26]. For more information on projection-depth-based medians, we refer the reader to [26, 27].
3 Main results
3.1 A private projection-depth-based median
In this section, we study a propose-test-release mechanism which outputs a private version of $\theta$, the projection-depth-based median. For common choices of $\mu,\ \sigma$, such as the median and median absolute deviation, the global sensitivity of the outlyingness function is not finite. Therefore, we cannot directly apply the exponential mechanism to develop a differentially private projection-depth-based median estimator. However, the median and median absolute deviation are robust statistics, in the sense that they are not corrupted under small levels of contamination. Thus, the outlyingness function should often be locally insensitive, making it natural to use in a PTR mechanism.
We now introduce the private projection-depth-based median. Let $W$ be a standard Laplace random variable. Take
$\lambda_{1}(\hat{\nu},\epsilon,\delta)=\Pr\left(\operatorname{SM}(\hat{\nu},\operatorname{EM}_{\hat{\nu},O})+\frac{2}{\epsilon}W>\frac{2\log(1/2\delta)}{\epsilon}\Bigg{|}\ \hat{\nu}\right),$ (3.1)
to be the test mechanism.
Definition 3.1.
Given parameters $\epsilon,\delta,\eta,\tau>0$, the private projection-depth-based median is $\tilde{\theta}\sim P_{\hat{\nu}}$, where $P_{\hat{\nu}}=P_{\hat{\nu}}(;\tau,\eta,\epsilon)$ is the PTR mechanism based on $(\operatorname{EM}_{\hat{\nu},O},\lambda_{1})$.
The private projection-depth-based median is $(\epsilon,\delta)$-differentially private (Theorem 5.1). We note that the algorithms of Brown et al., [6] and Liu et al., 2021b [18] are PTR algorithms based on exponential mechanisms and $\lambda_{1}$, using a different cost function. The performance of a PTR mechanism is naturally characterized by both the probability that the test is passed ($\mathbb{E}\lambda_{1}$) and the accuracy of the approximation of $\theta$ by $\tilde{\theta}$. With these two quantities, we characterize both the likelihood that the mechanism releases a useful statistic, and how useful that statistic is likely to be.
To analyze these quantities, we first introduce the projected sensitivity and the maximal absolute projected deviation of an arbitrary statistic, $T$. Define
$\tilde{S}_{n,k}(T)=\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k)\\ u\in\mathbb{S}^{d-1}\end{subarray}}|T(\hat{\nu}_{u})-T(\tilde{\nu}_{u})|\qquad\text{and}\qquad S_{n}(T)=\sup_{u\in\mathbb{S}^{d-1}}|T(\hat{\nu}_{u})-T(\nu_{u})|.$
The first quantity ($\tilde{S}_{n,k}(T)$), which we call the projected sensitivity, is an extension of sensitivity. It is the maximum change in $T$ when $k$ observations are set to arbitrary values, over all directions. The second quantity ($S_{n,k}(T)$), referred to as the maximal absolute projected deviation of $T$, is the maximum distance between the sample value of $T$ and the population value of $T$, over all directions. Next we introduce three regularity conditions.
Condition 1.
There exists constants $a_{1,d},a_{2,d,k}>0$, and a sequence $\kappa_{n}\longrightarrow 0$ as $n\rightarrow\infty$ such that for all $n,d\geq 1$ and $t\geq 0$ it holds that
$\Pr\left(\tilde{S}_{n,k}(T)\geq t\right)\leq a_{1,d}\exp\left(-a_{2,d,k}nt\right)+\kappa_{n}.$
Condition 2.
There exists a sequence of functions non-increasing functions $\zeta_{n,d}\longrightarrow 0$ as $n\to\infty$ such that for all $n,d\geq 1$ and $t\geq 0$ it holds that
$\Pr\Big{(}S_{n}(T)\geq t\Big{)}\lesssim\zeta_{n,d}(t).$
Condition 3.
There exists constants $c_{1},c_{2},c_{3}>0$ such that, for the population measure $\nu$, $0<c_{1}\leq\inf_{u\in\mathbb{S}^{d-1}}\sigma_{u}\leq\sup_{u\in\mathbb{S}^{d-1}}\sigma_{u}\leq c_{2}<\infty$ and $\sup_{u\in\mathbb{S}^{d-1}}|\theta^{\top}u-\mu_{u}|\leq c_{3}<\infty$.
Condition 1 can be viewed as guaranteeing robustness, with high probability. It says that the probability that one can contaminate $k$ observations in the sample so that the statistic of interest moves by at least $t$ is small. Of these three conditions, this first is the most difficult to check. As a result, we provide an example verifying it in Section 3.2. We show that for certain choices of $\mu$ and $\sigma$, Condition 1 reduces to requiring a lower bound on the “projected densities” of $\nu$. Condition 2 requires that $T(\hat{\nu}_{u})$ concentrates around its population value uniformly in $u$. This can typically be shown using standard empirical process or sub-Gaussian concentration techniques, (see for example, Lemma E.1). Condition 3 is standard for projection-depth-based estimators [27]. Under these three conditions, we derive a general bound on the probability that the proposed PTR mechanism fails the test based on $\lambda_{1}$.
Theorem 3.1.
Suppose that Conditions 13 hold for $(\nu,\mu,\sigma,k=4\log(1/2\delta)+1)$. For all $d,n\geq 1$, $\delta\leq 1/2$, $\epsilon,\eta,\tau>0$ such that $\tau>4\eta+16c_{3}/c_{1}$ and
$\delta>\exp\left(-\frac{\epsilon\tau}{8\eta}+\frac{\epsilon}{2}+d\log\left(\frac{2c_{2}(\tau+2\eta)+4c_{3}}{c_{1}\tau/4-\eta-4c_{3}}\right)\right),$
it holds that
$\mathbb{E}\lambda_{1}(\hat{\nu},\epsilon,\delta)\gtrsim 1-\delta-a_{1,d}\exp\left(-\frac{a_{2,d,4\log(1/2\delta)/\epsilon+1}c_{2}n(\eta\wedge 1)}{4(\tau\vee 1)}\right)-\kappa_{n}-\zeta_{n,,d}(c_{1}/4).$
The proof of Theorem 3.1 is given in Appendix B. To interpret Theorem 3.1 note that, given robust $\mu$ and $\sigma$ in the sense of Condition 1, the probability of failing the test in the mechanism is governed by the ratio $a_{2,d,4\log(1/2\delta)/\epsilon+1}\eta/\tau$ and $\delta$. Next, we show that the mechanism will be accurate with high probability. The following theorem bounds the deviation of the private projection-depth-based median around its population value.
Theorem 3.2.
Suppose Conditions 2 and 3 hold. Then, for all $\epsilon,\eta>0$, $\delta\leq 1/2$, $\gamma~{}\in~{}(3\zeta_{n,d}(c_{1}/2\wedge c_{3})\vee 3(1-\mathbb{E}\lambda_{1}(\hat{\nu},\epsilon,\delta)),1]$, $\tau\geq 20(2c_{3}/(10c_{2}-c_{1})+c_{3}/c_{1})$ and $n,d\geq 1$, with probability $1-\gamma$, it holds that
$\lVert\tilde{\theta}-\theta\rVert\lesssim\frac{c_{2}\eta}{\epsilon}\left[\log(1/\gamma)\vee d\log\left(\frac{2c_{2}\tau}{c_{3}}+1\right)\vee\frac{c_{3}}{c_{1}}\right]\vee\frac{c_{2}}{c_{1}}\zeta_{n,d}^{-1}(\gamma/3).$
The proof of Theorem 3.2 is given in Appendix B. Theorem 3.2 shows that, provided $\eta$ is relatively small and $\mathbb{E}\lambda_{1}$ is large enough, the cost of privacy for privately estimating the projection-depth-based median is governed by $\eta(\log(1/\gamma)\vee d)/\epsilon$. Theorem 3.1 and Theorem 3.2 quantify the trade-off inherent in $\eta$: smaller $\eta$ improves accuracy at the cost of a higher probability of failing the test. In general, one would like to show that the cost of privacy is less than the sampling error: $c_{2}\zeta_{n,d}^{-1}(\gamma/3)/c_{1}$. We will explore this in the next section.
3.2 Verifying Condition 1 and example applications
We next verify Condition 1 for the canonical choices of $\mu$ and $\sigma$, before applying Theorem 3.1 and Theorem 3.2 for Gaussian and Cauchy population measures. We demonstrate our methodology for the median and median absolute deviation, using the properties of order statistics.
For $\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$, let $F_{\nu}$ be the cumulative distribution function and $f_{\nu}$ be the Radon–Nikodym derivative of $\nu$. If $d=1$, let $F^{-1}_{\nu}$ be the left-continuous quantile function and let $\xi_{q,u}=F^{-1}_{\nu_{u}}(q)$. We require the following condition for the triple $(q,\nu,k)$, with $20<q\leq 1/$ 2.
Condition 4.
The measure $\nu$ is absolutely continuous and there exists $r,M>0$ such that, for the interval $\mathcal{I}=[\xi_{q-k/n,u}-r,\xi_{1-q+k/n,u}+r]$, we have $\inf_{\begin{subarray}{c}x\in\mathcal{I};u\in\mathbb{S}^{d-1}\end{subarray}}f_{\nu_{u}}(x)>M$.
For a given quantile, Condition 4 requires that the density of the projected population measure is in some interval which contains $(\xi_{q-k/n,u},\xi_{1-q+k/n,u})$, whose width depends on $k$. For $q=1/2$, Condition 4 is a multivariate extension of the condition required for the univariate PTR median [7], with a modified handling of $k$. Under Condition 4 the median and median absolute deviation both satisfy Conditions 1 and 2 (see Lemmas E.1 and E.2).
Theorem 3.3.
Suppose for a given $k\in[n]$, there exists $40<q^{\prime}<q<1/$ 4 such that both $k<n(q-q^{\prime})/2$ and Condition 4 holds for $(q^{\prime},\nu,k)$. Then the median and median absolute deviation satisfy Condition 1 with $a_{1,d}=\exp({d\log 9+\log(n(1-2q^{\prime}+1/n))})$, $a_{2,d,k}=3M/8k$ and $\kappa_{n}=\exp({-nr^{2}M^{2}+d\log(cn/d)}).$
The proof follows directly from Lemmas C.3 and C.4. Typically, we will have $k=o(n)$, and when we do, Theorem 3.3 is readily interpretable. Specifically, provided the (univariate) projected densities are bounded below when computed at the inner quantiles, then the median and median absolute deviation have small projected sensitivity with high probability. Note that we can prove an analogue of Theorem 3.3 with the interquartile range in place of the median absolute deviation (see Lemma C.5).
Remark 3.4.
Liu et al., 2021b [18] studied a projection-depth-based median taking $\mu$ and $\sigma$ to be the trimmed mean and variance, respectively. The projected sensitivities of trimmed estimators are governed by spacings between order statistics. The main technical argument in Theorem 3.3 is Lemma C.2, which controls the projected spacings between central order statistics. It is straightforward to apply these same techniques to show that Condition 1 holds for the trimmed estimators. Furthermore, the results of Fernandez-Rico, [13] can be applied to verify Condition 2, giving a straightforward application of Theorems 3.1 and 3.2.
To illustrate the methodology, we begin with a simple example, taking $\nu$ to be Gaussian.
Example 3.1 (Gaussian data).
Take the population measure to be $\mathcal{N}(\theta,\Sigma)$ with $\theta\in\mathbb{R}^{d}$ and $\Sigma$ as a positive definite symmetric matrix. Without loss of generality, we may assume that $\nu=\mathcal{N}(0,\Lambda)$, where $\Lambda$ is a diagonal matrix with eigenvalues $\beta_{1}\leq\cdots\leq\beta_{d}$ and let $\omega$ be the condition number of $\Lambda$, i.e., $\omega=\beta_{d}/\beta_{1}$. Specifically, the bounds produced by Theorems 3.1 and 3.2 when $\nu=\mathcal{N}(0,\Lambda)$ are equivalent to those produced by taking $\nu=\mathcal{N}(\theta,\Sigma)$, where $\Sigma=R\Lambda R^{\top}$ for some unitary matrix $R$. To see this, note that $R\nu+\theta=\mathcal{N}(\theta,\Sigma)$, where $A\nu+b$ represents the law of $AX+b$ if $X\sim\nu$. Given that the outlyingness function is invariant under affine transformations, with $O(x,\nu)=O(Ax+b,A\nu+b)$, and using the fact that the private median is computed using only the outlyingness function, the simplified case loses no generality.
First, we check Conditions 24. Direct calculation yields that Condition 3 is satisfied with $c_{1}\leq\beta_{1}^{1/2}$, $c_{2}\geq\beta_{d}^{1/2}$ and $c_{3}>0$. Take $c_{3}=1$. Next, we check Condition 4 for $q=1/8$, $k=4\log(1/2\delta)/\epsilon$, and $r=\beta_{1}^{1/2}/4$. Provided that $\delta>e^{-n\epsilon/32}/2$, then Condition 4 holds with
$M=(2\pi\beta_{d})^{-1/2}\exp({-(\Phi^{-1}(1/8-4\log(1/2\delta)/\epsilon n)-1/4)^{2}/2}).$
We now apply Theorem 3.3 to show that Condition 1 holds. If $4\log(1/2\delta)/\epsilon n\longrightarrow 0$, then $M\gtrsim\beta_{d}^{-1/2}$ for large $n$. Applying Theorem 3.3, $a_{1,d}\lesssim\exp({d+\log n})$, $a_{2,d,4\log(1/2\delta)/\epsilon}\gtrsim\epsilon(\beta_{d}^{1/2}\log(1/\delta))^{-1}$, and $\kappa_{n}\lesssim\exp\left(-c^{\prime}n/\omega+d\log(cd/n)\right).$ Next, using empirical process theory, (see Lemma E.1 and Lemma E.2) Condition 2 is satisfied for the median and median absolute deviation with
$\zeta_{n,d}(t)=\exp\left(-cnt^{2}/\beta_{d}+d\log(cn/d)\right)+\exp\left(-cn/\omega+d\log(cn/d)\right),$
for some universal constant $c>0$.
We can now apply Theorems 3.1 and 3.2. Beginning with Theorem 3.1, morally, for
$\delta\gtrsim\exp\left(-\frac{\epsilon\tau}{\eta}+\epsilon+d\log\omega\right)\vee e^{-n\epsilon},$
there exists universal constants $c,c^{\prime}>0$ such that
$\mathbb{E}\lambda(\hat{\nu},\epsilon,\delta)\gtrsim 1-\delta-\exp\left(-c^{\prime}\frac{n\eta\epsilon}{\log(1/\delta)(\tau\vee 1)}+d+\log n\right)-\exp\left(-\frac{c^{\prime}n}{\omega}+d\log(cn/d)\right).$
Thus, in order to control the probability of failing the test, it suffices to inspect the ratio $n\eta\epsilon/\log(1/\delta)\tau.$ Note that $O$ is a standardized measure of outlyingness, and doesn’t depend on the scale of the data, therefore it is natural to take $\tau$ to be a constant. Next, suppose $\delta=O(n^{-L})$ for some large $L>0$. Taking $\eta\gtrsim(L\log n)(d\vee L\log n)/n\epsilon$ results in $\mathbb{E}\lambda(\hat{\nu},\epsilon,\delta)\gtrsim 1-\delta$. In the best case, taking $L<<d$ results in the slowest scaling in the dimension, given by $\eta\gtrsim d\log^{2}n/n\epsilon$.
Next, taking $\eta\propto d\log^{2}n/n\epsilon$, we apply Theorem 3.2, which yields that, provided $n\geq d\omega$, for all $n^{-L}\lesssim\gamma\leq 1,$ and $d\geq 1$, with probability $1-\gamma$,
$\lVert\tilde{\theta}-\theta\rVert\lesssim\frac{d\beta_{d}^{1/2}\log^{2}n}{n\epsilon^{2}}\left[\log(1/\gamma)\vee d\log(\beta_{d}+1)\vee\beta_{1}^{-1/2}\right]\vee\sqrt{\beta_{d}\omega\frac{\log(1/\gamma)\vee d\log(n/d)}{n}}.$
For many reasonable values of $\gamma$, the sampling error greatly exceeds the cost of privacy. For instance, for $\gamma\propto\delta$, gives $||\tilde{\theta}-\theta||\approx\sqrt{\beta_{d}\omega\cdot d\log(n/d)/n}$ whenever, roughly, $n\gtrsim d^{3}\log^{4}n/\omega\epsilon^{4}$. Therefore, whenever $n\gtrsim d^{3}\log^{4}n/\omega\epsilon^{4}$, the deviations bound obtained from Theorem 3.2 matches the optimal deviations bound (for sub-Gaussian mean estimation) given by Cai et al., [8, see Theorem 3.1], see also [16], up to logarithmic terms and a factor of $\omega^{1/2}$. Of course, there exists estimators that achieve smaller upper bounds for private Gaussian mean estimation, e.g., [2, 6]. This is not surprising, given the usual robustness-efficiency trade-off in statistical estimation.
In our next example, we quantify the cost of privacy in the heavy-tailed setting, i.e., when no moments of $\nu$ exist. We show that even in this difficult setting, the cost of privacy is small relative to the sampling error. Furthermore, we observe that the “outlier error amplification” resulting from the heavy tails, greatly outweighs the cost of privacy.
Example 3.2 (Cauchy data).
Consider $\nu$ to be made up of $d$ independent Cauchy marginals with scale parameter equal to one and location parameter equal to zero. Analogous to Example 3.1, the assumptions of the scale parameter being one and location parameter being zero are made without loss of generality. Again, we check the Conditions 24. First, direct calculation yields that Condition 3 holds with $c_{1}=1$, $c_{2}=d^{1/2}$ and $c_{3}=1$.
Provided that $\delta\geq e^{-cn\epsilon}$ for some universal constant $c>0$, Condition 4 holds with $M\gtrsim d^{-1/2}$. Applying Theorem 3.3, we conclude that Condition 1 holds. Indeed, $a_{1,d}\lesssim\exp(d+\log n)$, $a_{2,d,4\log(1/2\delta)/\epsilon}\gtrsim\epsilon(d^{1/2}\log(1/\delta))^{-1}$, and $\kappa_{n}\lesssim\exp(-c^{\prime}n/d+d\log\left(cn/d\right))$. Next, we use Condition 4 to show that Condition 2 holds (see Lemma E.1 and Lemma E.2) with
$\zeta_{n,d}(t)=\exp\left(-c^{\prime}nt^{2}/d+d\log(cn/d)\right)+\exp\left(-c^{\prime}n/d^{2}+d\log(cn/d)\right),$
for universal constants $c,c^{\prime}>0$. We have now shown that all necessary conditions for the application of Theorems 3.1 and 3.2 are satisfied. We begin with an application of Theorem 3.1 and omit universal constants for clarity. For $\delta\gtrsim\exp\left(-\epsilon\tau/\eta+\epsilon+d\log(d\tau)\right)$, we have
$\mathbb{E}\lambda(\hat{\nu},\epsilon,\delta)$ $\gtrsim 1-\delta-\exp\left(-\frac{n\eta\epsilon}{d^{1/2}\log(1/\delta)\tau}+d+\log n\right)-\exp\left(-\frac{n}{d^{2}}+d\log(n/d)\right).$
Thus, compared to the Gaussian case, the lower bound on $\delta$ is larger. Furthermore, we require $\eta\gtrsim d^{3/2}\log^{2}n/n\epsilon$ to ensure that $\mathbb{E}\lambda(\hat{\nu},\epsilon,\delta)\gtrsim n^{-L}$, a factor of $d^{1/2}$ larger than in the Gaussian case. We can now apply Theorem 3.2. Taking $\eta\propto d^{3/2}\log^{2}n/n\epsilon$, for $n^{-L}<\gamma<1$ with probability $1-\gamma$, it holds that
$\lVert\tilde{\theta}-\theta\rVert\lesssim\frac{d^{3/2}\log^{2}n\left[\log(1/\gamma)\vee d\log d\right]}{n\epsilon^{2}}\vee\sqrt{\frac{d^{2}\log(1/\gamma)\vee d^{3}\log(n/d)}{n}}.$
It is immediate that the scaling in the dimension of the deviations bound is worse than that of the Gaussian case (Example 3.1). This is expected, given that heavy-tailed estimation is a more challenging problem. We refer to this as “outlier error amplification”. A similar phenomenon, in the form of sample complexity bounds, was observed by Ramsay et al., [20], for private medians based on other depth functions. Lastly, observe that the cost of privacy is still greatly outweighed by the sampling error, provided $n\gtrsim d^{2}\log^{4}n/\epsilon^{2}$.
4 Numerical experiments
To complement our theoretical results, we provide a numerical demonstration of our proposed methods. In order to compute the private projection-depth-based median, we employ discretized Langevin dynamics (as was done by Ramsay et al., [20]) and compute outlyingness by (randomly) discretizing the unit sphere, see Appendix A for the algorithm. More efficient algorithms could be constructed by using the methods and results of Dyckerhoff et al., [12], but we leave that for future work, as the current implementation’s differential privacy guarantee may be affected by the use of Markov chain Monte Carlo (MCMC) methods [22]. The purpose of the implementation is to numerically investigate the cost of privatizing the projection-depth-based median and compare to some other private measures of location, and the non-private projection-depth-based median. Our code is available on GitHub.
We compute the private and non-private projection-depth-based median based on the median, with the median absolute deviation as the scale estimate. We compare the performance of the projection-depth-based medians to the private smoothed integrated dual depth median [20], the non-private sample mean, and the private coin-press mean of Biswas et al., [5]. Following Ramsay et al., [20], for the integrated dual depth median, we set the smoothing parameter to 10, the prior as $\mathcal{N}(0,25dI)$. For the coin-press mean, we set the bounding ball to have a radius of $10\sqrt{d}$ and the algorithm was run for four iterations.
Figure 1: Empirical root mean squared error (ERMSE) of the location estimates under (left) Gaussian data and (right) contaminated Gaussian data. The cost of privacy is eclipsed by the outlier error amplification. The non-private estimates are all relatively close to their non-private counterparts. By contrast, the increase in error for all estimators resulting from contamination is apparent. The usual differences between robust and non-robust estimators are also apparent. Notably, the projection-depth-based median is the most robust.
Following Ramsay et al., [20], we simulated fifty instances of each PD median, with $n=10,000$ over a range of dimensions from two to twenty. The data were either generated from a standard Gaussian measure or contaminated standard Gaussian measure. In the contaminated model, 25% of the observations had mean $(5,\ldots,5)$. We set $\epsilon=10,\ \delta=10/n$, $\tau=1$ and $\eta=30(\log n)/n$. The number of vectors was chosen to be 500 for $d<20$ and 1000 otherwise.
Figure 1 displays the empirical root mean squared error (ERMSE) for the different location estimates. It is apparent that the cost of privacy is relatively low for all estimates, though, the classical trade-off between robustness and efficiency is evident. The mean estimators perform best in the uncontaminated setting. The opposite is true in the contaminated setting. The outlier error amplification is greater than the cost of privacy in both scenarios. That is, the increase in error in all estimates resulting from the outliers is much larger than the increase in error from private estimation. We also note that the private smoothed integrated dual depth median performs better than the private projection-depth-based median in the uncontaminated setting, but worse than the private projection-depth-based median in the contaminated setting. When considering which median to use, it is helpful to note that the smoothed integrated dual depth median is considerably faster to compute in high dimensions than the projection-depth-based median. However, it is only equivariant under similarity transformations [20], while the projection-depth-based medians, are equivariant under affine transformations [26].
5 General results on propose-test-release mechanisms
In this section, we elaborate on the notion of safety [6, 17] in the general case of PTR mechanisms. We show that safety is both sufficient and necessary for privacy of a PTR mechanism. We also present general results for analyzing a PTR mechanism based on the exponential mechanism. If $\lambda(\hat{\nu})$ is small whenever $\hat{\nu}$ is not in the safe set of $Q$ and the test mechanism is differentially private, then a PTR mechanism will be differentially private.
Theorem 5.1.
Given $\epsilon_{1},\epsilon_{2},\delta_{1},\delta_{2},\delta_{3}>0$, let $P$ be the PTR mechanism based on $Q(\cdot;\cdot,\epsilon_{2},\delta_{2})$ and $\lambda(\cdot,\epsilon_{1},\delta_{1})$. If $\lambda$ is $(\epsilon_{1},\delta_{1})$-differentially private and is such that for any $\hat{\nu}\notin\mathbf{S}(Q,\epsilon_{2},\delta_{2})$, $\lambda(\hat{\nu},\epsilon_{1},\delta_{1})<\delta_{3}$, then $P$ is $(\epsilon_{1}+\epsilon_{2},\delta_{1}+\delta_{2}\vee\delta_{3})$-differentially private.
Proof.
First, consider $\hat{\nu}\in\mathbf{S}(Q,\epsilon_{2},\delta_{2})$. By composition, the differential privacy bound with parameters $(\epsilon_{1}+\epsilon_{2},\delta_{1}+\delta_{2})$ will be satisfied. Now, consider $\hat{\nu}\notin\mathbf{S}(Q,\epsilon_{2},\delta_{3})$. For sets such that $\perp\notin A$, then $P_{\hat{\nu}}(A)\leq\lambda(\hat{\nu})\leq\delta_{3}.$ For sets such that $\perp\in A$, then
$P_{\hat{\nu}}(A)\leq P_{\hat{\nu}}(\{\perp\})+\delta_{3}\leq P_{\tilde{\nu}}(\{\perp\})e^{\epsilon_{1}}+\delta_{2}+\delta_{3}.\qed$
The previous theorem gives sufficient conditions on both the privacy guarantee of the test mechanism and the relationship between the test mechanism and the safe set of $Q$ to gain differential privacy of a PTR mechanism. We can conversely show that a $(\epsilon,\delta)$-differentially private propose test release algorithm consists of a test that is at least $(\epsilon,\delta)$-differentially private and, for datasets likely to pass the test, the release mechanism-dataset pair is safe.
Theorem 5.2.
Suppose that a PTR mechanism $P$ based on $Q$ and $\lambda$ is $(\epsilon,\delta)$-differentially private. Then
  1. 1.
    The test mechanism characterized by $\lambda$ satisfies $(\epsilon,\delta)$-differential privacy.
  2. 2.
    For any $\hat{\nu}$ such that $\lambda(\hat{\nu})>1/k$, the pair $(Q(;\hat{\nu},\epsilon),\hat{\nu})$ is $(2\epsilon,k\delta e^{\epsilon})$-safe.
Proof.
[…]
Now, the critical components for accuracy are that $\Pr\left(\hat{\nu}\in\mathbf{S}(Q,\epsilon,\delta)\right)$ is large and that $\lambda(\hat{\nu})$ is large when $\hat{\nu}\in\mathbf{S}(Q,\epsilon,\delta)$. We formalize this for the well-known exponential mechanism as follows. Let $G$ be the CDF of a standard Laplace RV.
Lemma 5.3.
For all $\epsilon,\tau,\eta>0$, $\delta\leq e^{-\epsilon/2}/2$, $n,d\geq 1$ and $\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$ it holds that
$\mathbb{E}\lambda(\hat{\nu},\epsilon,\delta)\geq 1-\Pr\left(\inf_{y>0}e^{-\epsilon y/4\eta}\frac{\operatorname{vol}(A_{\tau+2\eta,\hat{\nu}}\cap A_{\tau-2\eta,\hat{\nu}}^{c})}{\operatorname{vol}(A_{\tau-y-2\eta,\hat{\nu}})}>e^{-\epsilon/4}\delta/3\right)\\ -\inf_{k\in\{0\}\cup[n-1]}\left[\Pr\left(\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k+1)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|\phi(x,\hat{\nu})-\phi(x,\tilde{\nu})|>\eta/2\right)+G(\log(1/2\delta)-\epsilon k/2)\right].$
The proof of Lemma 5.3 can be found in Appendix D. This lemma can be interpreted as follows: the probability of passing the test is high when the probability that, for some small $\alpha>0$, the $\tau\pm\alpha$-contours of $\phi(\cdot,\hat{\nu})$ are not too steep and the objective is locally insensitive within the $\tau\pm\alpha$-contour.
6 Conclusion
In this paper we introduce $(\epsilon,\delta)$-differentially private projection-depth-based medians. Without enforcing any moment bounds on the population measure, we bound the deviation of the private projection-depth-based medians about their population values, and quantify the probability that the test in the PTR algorithm fails. We demonstrate our results by applying them to the particular projection-depth-based median that uses the median and median absolute deviation. In this case, we show that in the Gaussian setting, the deviation bound nearly matches the lower bound for sub-Gaussian mean estimation given by Cai et al., [8]. This provides yet another example of the usefulness of robust statistics in a differentially private setting. We also show that in the Cauchy setting, the “outlier error amplification” is larger than the cost of privacy. This phenomenon is also observed in our numerical studies.
The main concern for future work is the computation of private projection-depth-based medians. In particular, it is interesting and important to quantify the violation of the differential privacy guarantee, if any, that results from using Langevin Dynamics to sample from the exponential mechanism. In addition, more sophisticated methods to compute the outlyingness could be developed, possibly using the results of Dyckerhoff et al., [12].
Appendix A Algorithm
Here we outline the algorithm used to compute the private, projection-depth-based median in Section 4. We will approximate $O(x,\hat{\nu})$ by $\widehat{O}(x,\hat{\nu})=\sup_{u\in\mathcal{U}_{N}}O_{u}(x,\hat{\nu})$, where $\mathcal{U}_{N}$ is a set of $N$ unit vectors sampled uniformly from $\mathbb{S}^{d-1}$. The algorithm consists of two parts, executing the test and sampling from the exponential mechanism. For the latter, observe that the outlyingness function is quasiconvex, i.e., its lower level sets are convex. This allows us to apply an envelope theorem to compute the gradient of $\widehat{O}_{u}(x,\hat{\nu})$ with respect to $x$. To this end, for $x\in A_{\tau}$, let $u_{x}$ be the unit vector that maximizes $\widehat{O}_{u}(x,\hat{\nu})$. Then, for any $x\in A_{\tau}$, it holds that
$\nabla\widehat{O}(x,\nu)$ $=u_{x}\frac{\mathrm{sign}(x^{\top}u_{x}-\mu_{u_{x}})}{\sigma_{u_{x}}}.$
To compute the non-private projection-depth-based median, we use gradient descent, and, to sample from the exponential mechanism, we use Langevin Dynamics. Given that Langevin Dynamics samples approximately from the exponential mechanism, we note that this may violate the differential privacy condition, and stress that this implementation is just to confirm the theoretical properties of the proposed private median. Of course, more sophisticated methods to compute the outlyingness could be developed, such as those based on [12], but we leave that for future work. Indeed, given the caveat on using Langevin Dynamics to sample from the exponential mechanism mentioned above, an algorithm to compute the private projection-depth-based median is an interesting open problem.
Algorithm 1 Exponential Mechanism via Langevin Dynamics
procedure EM($X_{1},\ldots,X_{n},\epsilon,\eta,\tau,N,T,\omega$)
     Set $\mathcal{U}_{N}$ to be $N$ independent random vectors uniformly over $\mathbb{S}^{d-1}$.
     Set $Y_{0}=\mathcal{N}(0,I)$
     for $i\in[T]$ do
         Set $O=\widehat{O}(x,\hat{\nu})$
         if $O>\tau$ then
              Set $\nabla O=0$
         else
              Set $\nabla O=\nabla\widehat{O}(x,\hat{\nu})$
         end if
         Set $Z_{i}=\mathcal{N}(0,I)$
         Set $Y_{i}=Y_{i-1}-\omega\frac{\epsilon}{4\eta}\nabla O+\sqrt{2\omega}Z_{i}$
     end for
return $Y_{T}$
end procedure
Algorithm 2 Test Mechanism
procedure Test($X_{1},\ldots,X_{n},\epsilon,\delta,\eta,\tau,d$)
     Set $k=1$
     Set $\mathrm{Dev\_Cond}=$True
     Set $\mathrm{Vol\_Cond}=$True
     while $\mathrm{Dev\_Cond}$ and $\mathrm{Vol\_Cond}$ do
         Set $k=k+1$
         Set $\mathrm{Vol\_Cond}=\mathrm{VR}_{k}<\delta$
         Compute $\widehat{S}_{n,k}(\sigma),\ \widehat{S}_{n,k}(\mu),\ \hat{b}$
         Set $\mathrm{Dev\_Cond}=\widehat{\Delta}_{k}<\eta$
     end while
     Set $\operatorname{SM}=k-2$.
     Generate $W\sim\operatorname*{{\rm Laplace}}(0,1)$
return $\operatorname{SM}+\frac{2}{\epsilon}>\frac{2\log(1/2\delta)}{\epsilon}$
end procedure
Algorithm 3 Private PDB Median
procedure Private_Median($X_{1},\ldots,X_{n},\epsilon,\delta,\eta,\tau,d,T,\omega$)
     Set $\mathrm{test}=$Test($X_{1},\ldots,X_{n},\epsilon,\delta,\eta,\tau,d$)
     if $\mathrm{test}$ then return $\perp$
     else  return $\operatorname{EM}(X_{1},\ldots,X_{n},\epsilon,\eta,\tau,T,\omega)$
     end if
end procedure
To perform the test, it is clearly difficult to compute the safety margin in this case. As a result, we can make use of Lemma D.2, Lemma B.3 and Lemma B.4. Now, define
$\widehat{S}_{n,k}(T)=\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k)\\ u\in\mathcal{U}_{N}\end{subarray}}|T(\hat{\nu}_{u})-T(\tilde{\nu}_{u})|.$
We must first compute three quantities: $\widehat{S}_{n,k}(\mu),\widehat{S}_{n,k}(\sigma)$ and $\hat{b}=\inf_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k)\cup\{\hat{\nu}\}\\ u\in\mathcal{U}_{N}\end{subarray}}\sigma(\hat{\nu}_{u})$. First, it is easy to check that if $\mu=\operatorname{MED}$, then
$\widehat{S}_{n,k}(\mu)\leq\sup_{u\in\mathcal{U}_{N}}[|(X^{\top}u)_{(\lfloor(n+1)/2\rfloor+k)}-(X^{\top}u)_{(\lfloor(n+1)/2\rfloor)}|\vee|(X^{\top}u)_{(\lfloor(n+1)/2\rfloor)}-(X^{\top}u)_{(\lfloor(n+1)/2\rfloor-k)}|].$
In order to compute the $\widehat{S}_{n,k}(\sigma)$ where $\sigma=\operatorname{MAD}$, we can first consider the local sensitivity of the median absolute deviation, for a univariate sample $Z_{1},\ldots,Z_{n}$ with empirical measure $\hat{\nu}$. Let $Y_{i}(z)=|Z_{i}-z|$ and call this measure $\hat{\nu}_{z}$. Consider $\tilde{\nu}\in\mathcal{M}(k,\hat{\nu})$. First, note that $\mu(\tilde{\nu})\in[Z_{(\lfloor(n+1)/2\rfloor+k)},Z_{(\lfloor(n+1)/2\rfloor-k)}]=[a_{k},b_{k}]$. Now observe that
$\sup_{\tilde{\nu}\in\mathcal{M}(k,\hat{\nu})}|\sigma(\tilde{\nu})-\sigma(\hat{\nu})|$ $\leq\sup_{z\in[a_{k},b_{k}]}\sup_{\tilde{\nu}\in\mathcal{M}(k,\hat{\nu}_{z})}|\operatorname{MED}(\hat{\nu}_{z})-\operatorname{MED}(\tilde{\nu})|$
$\leq\sup_{z\in[a_{k},b_{k}]}|Y_{(\lfloor(n+1)/2\rfloor+k)}(z)-Y_{(\lfloor(n+1)/2\rfloor)}(z)|$
$\hskip 100.0pt\vee|Y_{(\lfloor(n+1)/2\rfloor-k)}(z)-Y_{(\lfloor(n+1)/2\rfloor)}(z)|.$
Note that the right-hand side of the above inequality is a piecewise linear function in $z$, with knots at the observations. Thus, it suffices to compute this function for the $2k$ observations in $[a_{k},b_{k}]$. Thus, letting $Y_{i,u}(z)=|X_{i}^{\top}u-z|$,
$\widehat{S}_{n,k}(\sigma)\leq\sup_{\begin{subarray}{c}u\in\mathcal{U}_{N}\\ z\in[a_{k},b_{k}]\end{subarray}}|Y_{(\lfloor(n+1)/2\rfloor+k),u}(z)-Y_{(\lfloor(n+1)/2\rfloor),u}(z)|\vee|Y_{(\lfloor(n+1)/2\rfloor-k),u}(z)-Y_{(\lfloor(n+1)/2\rfloor),u}(z)|.$
Similarly, $\hat{b}\leq\inf_{u\in\mathcal{U}_{N}}\inf_{z\in[a_{k},b_{k}]}|Y_{(\lfloor(n+1)/2\rfloor),u}(z)|.$
Next, we can compute $\widehat{\Delta}_{k}=((\tau+\eta)\widehat{S}_{n,k}(\sigma)+\widehat{S}_{n,k}(\mu))/\hat{b}$. The last step is to compute the volume condition in Lemma D.2 using Lemma B.4 and check if
$\mathrm{VR}_{k}\coloneqq\exp\left(-\epsilon\tau/4\eta+\epsilon(1/2+\hat{\alpha}_{\mu}/4\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}\eta)+d\right)\left(\frac{\sup_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau+2\eta)+\hat{\alpha}_{\mu}}{\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}4\eta d\epsilon}\right)^{d}\leq\delta.$
All the described methods are summarized in Algorithms 13.
Appendix B Proofs from Section 3.1
Prior to proving the main results, we first establish probability bounds that will be used throughout the following proofs.
Lemma B.1.
Let $\hat{\alpha}_{\mu}=\sup_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}-\inf_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}$ and $\alpha_{\mu}=\sup_{u\in\mathbb{S}^{d-1}}\mu_{u}-\inf_{u\in\mathbb{S}^{d-1}}\mu_{u}$.
  1. 1.
    Assuming Conditions 1-3:
    $\Pr\Big{(}\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}\sigma(\tilde{\nu})\leq\frac{c_{1}}{2}\Big{)}\lesssim a_{1,d}e^{-a_{2,d,k}nc_{1}/4}+\kappa_{n}+\zeta_{n,d}(\frac{c_{1}}{4}).$ (B.1)
  2. 2.
    Assuming $\alpha_{\mu}\leq 2c_{3}$ and that Condition 2 holds:
    $\Pr\left(\widehat{a}_{\mu}>4c_{3}\right)\lesssim\zeta_{n,d}(c_{3}).$ (B.2)
  3. 3.
    Assuming Conditions 2 and 3:
    $\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})\geq 2c_{2}\right)\lesssim\zeta_{n,d}(c_{2}).$ (B.3)
  4. 4.
    Assuming Conditions 2 and 3:
    $\Pr\left(\inf_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})\leq c_{1}/2\right)\lesssim\zeta_{n,d}(\frac{c_{1}}{2}).$ (B.4)
Proof.
Proof of (B.1):
$\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}\sigma(\tilde{\nu})$ $=\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}(\sigma(\tilde{\nu})-\sigma(\hat{\nu})+\sigma(\hat{\nu}))$
$\geq\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}(\sigma(\tilde{\nu})-\sigma(\hat{\nu}))+\inf_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu})$
$\geq\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}(\sigma(\tilde{\nu})-\sigma(\hat{\nu}))+\inf_{u\in\mathbb{S}^{d-1}}(\sigma(\hat{\nu})-\sigma(\nu))+\inf_{u\in\mathbb{S}^{d-1}}\sigma(\nu)$
$\geq-\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}|\sigma(\tilde{\nu})-\sigma(\hat{\nu})|-\sup_{u\in\mathbb{S}^{d-1}}|\sigma(\hat{\nu})-\sigma(\nu)|+c_{1}.$
The last inequality uses Condition 3. Then, if $\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}\sigma(\tilde{\nu})\leq c_{1}/2$,
$-\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}|\sigma(\tilde{\nu})-\sigma(\hat{\nu})|-\sup_{u\in\mathbb{S}^{d-1}}|\sigma(\hat{\nu})-\sigma(\nu)|+c_{1}$ $\leq\frac{c_{1}}{2}$
$\implies\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}|\sigma(\tilde{\nu})-\sigma(\hat{\nu})|+\sup_{u\in\mathbb{S}^{d-1}}|\sigma(\hat{\nu})-\sigma(\nu)|$ $\geq\frac{c_{1}}{2}.$
Taken together,
$\Pr\left(E_{2}^{c}\right)$ $\leq\Pr\left(\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}|\sigma(\tilde{\nu})-\sigma(\hat{\nu})|+\sup_{u\in\mathbb{S}^{d-1}}|\sigma(\hat{\nu})-\sigma(\nu)|\geq\frac{c_{1}}{2}\right)$
$\leq\Pr\left(\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}|\sigma(\tilde{\nu})-\sigma(\hat{\nu})|\geq\frac{c_{1}}{4}\right)+\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}|\sigma(\hat{\nu})-\sigma(\nu)|\geq\frac{c_{1}}{4}\right)$
$\lesssim a_{1,d}e^{-a_{2,d,k}nc_{1}/4}+\kappa_{n}+\zeta_{n,d}(\frac{c_{1}}{4}).$
The final inequality follows directly from Conditions 1 and 2
Proof of (B.2): Note that
$\hat{\alpha}_{\mu}=\sup_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}-\inf_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}$ $=\sup_{u\in\mathbb{S}^{d-1}}(\hat{\mu}_{u}+\mu_{u}-\mu_{u})+\sup_{u\in\mathbb{S}^{d-1}}(-\hat{\mu}_{u}+\mu_{u}-\mu_{u})$
$\leq\sup_{u\in\mathbb{S}^{d-1}}|\hat{\mu}_{u}-\mu_{u}|+\sup_{u\in\mathbb{S}^{d-1}}\mu_{u}+\sup_{u\in\mathbb{S}^{d-1}}|\mu_{u}-\hat{\mu}_{u}|-\inf_{u\in\mathbb{S}^{d-1}}\mu_{u}$
$=2\sup_{u\in\mathbb{S}^{d-1}}|\mu_{u}-\hat{\mu}_{u}|+a_{\mu}$
$\leq 2\left(\sup_{u\in\mathbb{S}^{d-1}}|\mu_{u}-\hat{\mu}_{u}|+c_{3}\right),$
where we used the fact that $\alpha_{\mu}\leq 2c_{3}$. As a result, applying Condition 2,
$\Pr\left(E_{3}^{c}\right)\leq\Pr\left(2\left(\sup_{u\in\mathbb{S}^{d-1}}|\mu_{u}-\hat{\mu}_{u}|+c_{3}\right)>4c_{3}\right)=\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}|\mu_{u}-\hat{\mu}_{u}|>c_{3}\right)\lesssim\zeta_{n,d}(c_{3}).$
Proof of (B.3):
$\sup_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})$ $=\sup_{u\in\mathbb{S}^{d-1}}\left(\sigma(\hat{\nu}_{u})-\sigma(\nu_{u})+\sigma(\nu_{u})\right)$
$\leq\sup_{u\in\mathbb{S}^{d-1}}\left(\sigma(\hat{\nu}_{u})-\sigma(\nu_{u})\right)+\sup_{u\in\mathbb{S}^{d-1}}\sigma(\nu_{u})$
$\leq\sup_{u\in\mathbb{S}^{d-1}}|\sigma(\hat{\nu}_{u})-\sigma(\nu_{u})|+c_{2},$
where the final inequality relies on condition 3. Then,
$\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})\geq 2c_{2}\right)$ $\leq\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}|\sigma(\hat{\nu}_{u})-\sigma(\nu_{u})|+c_{2}\geq 2c_{2}\right)$
$=\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}|\sigma(\hat{\nu}_{u})-\sigma(\nu_{u})|\geq c_{2}\right)$
$\lesssim\zeta_{n,d}(c_{2}),$
where the final result is a straightforward application of condition 2.
Proof of (B.4):
$\inf_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})$ $=\inf_{u\in\mathbb{S}^{d-1}}\left(\sigma(\hat{\nu}_{u})-\sigma(\nu_{u})+\sigma(\nu_{u})\right)$
$\geq\inf_{u\in\mathbb{S}^{d-1}}\left(\sigma(\hat{\nu}_{u})-\sigma(\nu_{u})\right)+\inf_{u\in\mathbb{S}^{d-1}}\left(\sigma(\nu_{u})\right)$
$=-\sup_{u\in\mathbb{S}^{d-1}}\left(\sigma(\nu_{u})-\sigma(\hat{\nu}_{u})\right)+\inf_{u\in\mathbb{S}^{d-1}}\left(\sigma(\nu_{u})\right)$
$\geq-\sup_{u\in\mathbb{S}^{d-1}}\left|\sigma(\nu_{u})-\sigma(\hat{\nu}_{u})\right|+c_{1},$
where the final inequality relies on Condition 3. This means that
$\Pr\left(\inf_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})\leq\frac{c_{1}}{2}\right)$ $\leq\Pr\left(-\sup_{u\in\mathbb{S}^{d-1}}\left|\sigma(\nu_{u})-\sigma(\hat{\nu}_{u})\right|+c_{1}\leq\frac{c_{1}}{2}\right)$
$=\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}\left|\sigma(\nu_{u})-\sigma(\hat{\nu}_{u})\right|\geq\frac{c_{1}}{2}\right)$
$\lesssim\zeta_{n,d}(\frac{c_{1}}{2}),$
where the final line is a simple application of Condition 2. ∎
Proof of Theorem 3.1.
First, note that $\mu$ and $\sigma$ may satisfy Conditions 1 and 2 for different parameters, e.g., different sequences $\zeta$. However, we can define a set of joint parameters $\zeta,\kappa_{n},\ldots$ for which the bounds in Conditions 1 and 2 hold for both $\mu$ and $\sigma$, which we will use throughout the proof. Let $b_{k,u}^{-}=\inf_{\tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)}\sigma(\tilde{\nu})$, $\hat{\alpha}_{\mu}=\sup_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}-\inf_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}$, $\alpha_{\mu}=\sup_{u\in\mathbb{S}^{d-1}}\mu_{u}-\inf_{u\in\mathbb{S}^{d-1}}\mu_{u}$.
An application of Lemma 5.3 with $\phi=O$, and subsequently taking $k=4\log(1/2\delta)/\epsilon$ yields:
$\mathbb{E}\lambda(\hat{\nu},\epsilon,\delta)$ $\geq 1-\Pr\left(E_{1}\right)$
$\hskip 30.0pt-\inf_{k\in[n]}\left[\Pr\Big{(}\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k+1)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|O(x,\hat{\nu})-O(x,\tilde{\nu})|>\eta\Big{)}+G\left(\log\frac{1}{2\delta}-\frac{\epsilon k}{2}\right)\right]$ (B.7)
$\geq 1-\Pr\left(E_{1}\right)-\Pr\Big{(}\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},4\log(1/2\delta)/\epsilon+1)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|O(x,\hat{\nu})-O(x,\tilde{\nu})|>\eta\Big{)}-\delta,$ (B.10)
where
$E_{1}=\left\{\inf_{y>0}e^{-\epsilon y/4\eta}\frac{\operatorname{vol}(A_{\tau+2\eta,\hat{\nu}}\cap A^{c}_{\tau-2\eta,\hat{\nu}})}{\operatorname{vol}(A_{\tau-y-2\eta,\hat{\nu}})}>e^{-\epsilon/2}\delta\right\}.$
We now bound the deviation of the outlyingness function with high probability. To this end, note that Lemma B.3 implies that for any $k\in[n]$,
$\Pr\Big{(}\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|O(x,\hat{\nu})-O(x,\tilde{\nu})|>\eta\Big{)}\leq\Pr\left(\Delta_{k}>\eta\right),$
where
$\Delta_{k}=\frac{\tau\tilde{S}_{n,k}(\sigma)+\tilde{S}_{n,k}(\mu)}{\inf_{u\in\mathbb{S}^{d-1}}b_{k,u}^{-}}.$
Define $E_{2}=\left\{\inf_{u\in\mathbb{S}^{d-1}}b_{k,u}^{-}\geq c_{1}/2\right\}$. Then, conditioning on $E_{2}$ we have that
$\Delta_{k}\leq\frac{2}{c_{1}}\left[\tau\tilde{S}_{n,k}(\sigma)+\tilde{S}_{n,k}(\mu)\right],$
and so
$\Pr\left(\Delta_{k}\geq\eta|E_{2}\right)$ $\leq\Pr\left(\frac{2}{c_{1}}\left[\tau\tilde{S}_{n,k}(\sigma)+\tilde{S}_{n,k}(\mu)\right]\geq\eta|E_{2}\right)$
$\leq\Pr\left(\frac{2}{c_{1}}\left[\tau\tilde{S}_{n,k}\right]\geq\eta/2|E_{2}\right)+\Pr\left(\frac{2}{c_{1}}\left[(\sigma)+\tilde{S}_{n,k}(\mu)\right]\geq\eta/2|E_{2}\right)$
$=\Pr\left(\tilde{S}_{n,k}\geq c_{1}\eta/4\tau|E_{2}\right)+\Pr\left((\sigma)+\tilde{S}_{n,k}(\mu)\geq c_{1}\eta/4|E_{2}\right).$
Thus, applying Condition 1 gives
$\Pr\left(\Delta_{k}>\eta|E_{2}\right)$ $\lesssim\Pr\left(\tilde{S}_{n,k}(\mu)>c_{1}\eta/4|E_{2}\right)+\Pr\left(\tilde{S}_{n,k}(\sigma)>c_{1}\eta/4\tau|E_{2}\right)$
$\lesssim a_{1,d}e^{-a_{2,d,k}c_{1}n\eta/4(\tau\vee 1)}+\kappa_{n}.$ (B.11)
Note $\Pr\left(E_{2}^{c}\right)$ is bounded based on (B.1). Combining these past three results yields
$\Pr\Big{(}\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|O(x,\hat{\nu})-O(x,\tilde{\nu})|>\eta\Big{)}$
$\qquad\leq\Pr\left(\Delta_{k}>\eta|E_{2}\right)\Pr\left(E_{2}\right)+\Pr\left(\Delta_{k}>\eta|E_{2}^{c}\right)\Pr\left(E_{2}^{c}\right)$
$\qquad\leq\Pr\left(\Delta_{k}>\eta|E_{2}\right)+\Pr\left(E_{2}^{c}\right)$
$\qquad\lesssim a_{1,d}\exp\left(-a_{2,d,k}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n,d}(\frac{c_{1}}{4}).$
We next bound $\Pr\left(E_{1}\right)$. To do so we will make use of Lemma B.4. In order to apply the result, we require that $\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-y-2\eta)-\hat{\alpha}_{\mu}>0$. Note that, given $E_{2}$ and the restriction that $\tau\geq 4\eta+16c_{3}/c_{1}$,
$\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-y-2\eta)-\hat{\alpha}_{\mu}\geq\frac{c_{1}(\tau-y-2\eta)}{2}-\hat{\alpha}_{\mu}\geq c_{1}\eta+8c_{3}-\frac{c_{1}}{2}y-\hat{\alpha}_{\mu}.$
If we take $E_{3}=\left\{\hat{\alpha}_{\mu}\leq 4c_{3}\right\}$, then further conditioning on $E_{3}$ gives that, for all $y\in\mathcal{L}$, where $\mathcal{L}=(0,\tau-2\eta-8c_{3}/c_{1})$, $\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-y-2\eta)-\hat{\alpha}_{\mu}>0$. Thus,
$\frac{\operatorname{vol}(A_{\tau+2\eta,\hat{\nu}}\cap A^{c}_{\tau-2\eta,\hat{\nu}})}{\operatorname{vol}(A_{\tau-y-2\eta,\hat{\nu}})}$ $\leq\left(\frac{\sup_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau+2\eta)+\hat{\alpha}_{\mu}}{\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-y-2\eta)-\hat{\alpha}_{\mu}}\right)^{d}$
$\qquad-\left(\frac{\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-2\eta)-\hat{\alpha}_{\mu}}{\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-y-2\eta)-\hat{\alpha}_{\mu}}\right)^{d}$
$\leq\left(\frac{\sup_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau+2\eta)+\hat{\alpha}_{\mu}}{\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-y-2\eta)-\hat{\alpha}_{\mu}}\right)^{d}.$
Define $E_{4}=\left\{\sup_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})\leq 2c_{2}\right\}$, so that given $E_{5}=E_{2}\cap E_{3}\cap E_{4}$,
$\left(\frac{\sup_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau+2\eta)+\hat{\alpha}_{\mu}}{\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-y-2\eta)-\hat{\alpha}_{\mu}}\right)^{d}\leq\left(\frac{2c_{2}(\tau+2\eta)+4c_{3}}{c_{1}(\tau-y-2\eta)/2-4c_{3}}\right)^{d}.$
Thus, taken together,
$\Pr\left(E_{1}\right)$ $\leq\Pr\left(E_{1}|E_{5}\right)+\Pr\left(E_{5}^{c}\right)$
$\leq\Pr\left(\inf_{y\in\mathcal{L}}e^{-\epsilon y/4\eta}\left(\frac{2c_{2}(\tau+2\eta)+4c_{3}}{c_{1}(\tau-y-2\eta)/2-4c_{3}}\right)^{d}>e^{-\epsilon/2}\delta\right)+\Pr\left(E_{5}^{c}\right)$
$\leq\mathbbm{1}\left\{\inf_{y\in\mathcal{L}}e^{\epsilon/2-\epsilon y/4\eta}\left(\frac{2c_{2}(\tau+2\eta)+4c_{3}}{c_{1}(\tau-y-2\eta)/2-4c_{3}}\right)^{d}>\delta\right\}+\Pr\left(E_{5}^{c}\right).$
Taking $y=\tau/2$, the inequality in the indicator is never satisfied for any
$\delta>e^{\epsilon/2-\epsilon y/4\eta}\left(\frac{2c_{2}(\tau+2\eta)+4c_{3}}{c_{1}(\tau-y-2\eta)/2-4c_{3}}\right)^{d}=\exp\left\{\frac{\epsilon}{2}-\frac{\epsilon\tau}{8\eta}+d\log\left(\frac{2c_{2}(\tau+2\eta)+4c_{3}}{c_{1}\tau/4-c_{1}\eta-4c_{3}}\right)\right\}.$
Thus, for any such $\delta$, we have $\Pr\left(E_{1}\right)\leq\Pr\left(E_{5}^{c}\right)$.
Using equations (B.1), (B.2), and (B.3) we have bounds on $\Pr\left(E_{2}^{C}\right)$, $\Pr\left(E_{3}^{C}\right)$, and $\Pr\left(E_{4}^{C}\right)$ respectively. These three bounds combine to give
$\Pr\left(E_{5}^{c}\right)\lesssim a_{1,d}e^{-a_{2,d,k}nc_{1}/4}+\kappa_{n}+\zeta_{n,d}(\frac{c_{1}}{4}\wedge c_{3}).$
Finally,
$\mathbb{E}\lambda(\hat{\nu},\epsilon,\delta)$ $\geq 1-\Pr\left(E_{1}\right)-\Pr\Big{(}\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},4\log(1/2\delta)/\epsilon+1)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|O(x,\hat{\nu})-O(x,\tilde{\nu})|>\eta\Big{)}-\delta$
$\gtrsim 1-\Pr\left(E_{5}^{c}\right)-\left(a_{1,d}\exp\left(-a_{2,d,k}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n,d}(\frac{c_{1}}{4})\right)-\delta$
$\gtrsim 1-(a_{1,d}e^{-a_{2,d,k}nc_{1}/4}+\kappa_{n}+\zeta_{n,d}(\frac{c_{1}}{4}\wedge c_{3}))$
$\qquad-\left(a_{1,d}\exp\left(-a_{2,d,k}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n,d}(\frac{c_{1}}{4})\right)-\delta$
$\gtrsim 1-\exp\left(-a_{2,d,k}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)-\kappa_{n}-\zeta_{n,d}(\frac{c_{1}}{4})-\delta\qed.$
Proof of Theorem 3.2.
To prove Theorem 3.2, it suffices to find a lower bound on $t$ such that
$\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t\right)\leq\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t\big{|}\tilde{\theta}\neq\perp\right)+\Pr\left(\tilde{\theta}=\perp\right)\leq\gamma.$
Furthermore, by assumption, $\Pr\left(\tilde{\theta}=\perp\right)=1-\mathbb{E}\lambda(\hat{\nu},\epsilon,\delta)\leq\gamma/3$. Thus, it suffices to find $t$ such that
$\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t\big{|}\tilde{\theta}\neq\perp\right)\leq 2\gamma/3.$
To this end, applying Lemma B.2, yields
$\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t\big{|}\tilde{\theta}\neq\perp\right)\\ \lesssim\inf_{\begin{subarray}{c}0<s<c_{1}\tau/2-c_{3}\\ 0<r<c_{1}\tau/2-s-c_{3}\end{subarray}}\Bigg{[}\exp\left(\frac{\epsilon(c_{1}/2+2c_{2})(c_{3}+s)}{4\eta c_{1}c_{2}}-\frac{\epsilon t}{8\eta c_{2}}+\frac{\epsilon r}{2c_{1}\eta}+\log\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}\right)\\ +\zeta_{n,d}\left(s\wedge c_{3}\wedge c_{1}/2\right)\Bigg{]}.$ (B.12)
We first show that
$\inf_{\begin{subarray}{c}0<s<\frac{c_{1}\tau}{2}-c_{3}\\ 0<r<\frac{c_{1}\tau}{2}-s-c_{3}\end{subarray}}\Bigg{[}\exp\left(\frac{\epsilon(\frac{c_{1}}{2}+2c_{2})(c_{3}+s)}{4\eta c_{1}c_{2}}-\frac{\epsilon t}{8\eta c_{2}}+\frac{\epsilon r}{2c_{1}\eta}+\log\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}\right)\Bigg{]}\leq\frac{\gamma}{3}.$ (B.13)
To this end, using the fact that $c_{1}<c_{2}$, for any $s,r>0$, we have that
$\frac{(c_{1}/2+2c_{2})(c_{3}+s)}{4c_{1}c_{2}}-\frac{t}{8c_{2}}+\frac{r}{2c_{1}}$ $\leq\frac{5(c_{3}+s)}{8c_{1}}-\frac{t}{8c_{2}}+\frac{r}{2c_{1}}=\frac{5(c_{3}+s)+4r-tc_{1}/c_{2}}{8c_{1}}.$
Next, for some $k_{r},k_{s}>0$, take $r=tc_{1}/(k_{r}c_{2})$ and $s=r/k_{s}$. Then
$\frac{5(c_{3}+s)+4r-tc_{1}/c_{2}}{8c_{1}}$ $=\frac{(5+4k_{s})tc_{1}/(k_{r}c_{2}k_{s})-tc_{1}/c_{2}+5c_{3}}{8c_{1}}=\frac{(5+4k_{s}-k_{r}k_{s})t}{8c_{2}k_{r}k_{s}}+\frac{5c_{3}}{8c_{1}}.$
In order to apply this result to (B.13), it must be the case that $0<s<c_{1}\tau/2-c_{3}$ and $0<r<c_{1}\tau/2-s-c_{3}$. Note that $r<c_{1}\tau/2-s-c_{3}\implies r+s<c_{1}\tau/2-c_{3}\implies s<c_{1}\tau/2-c_{3}$, so it suffices to consider $s>0$, and $0<r<c_{1}\tau/2-s-c_{3}$. First, observe that $s,r>0$ always holds. Next, simplifying the second bound yields
$r<c_{1}\tau/2-s-c_{3}$ $\implies r+s<c_{1}\tau/2-c_{3}$
$\implies\frac{tc_{1}}{k_{r}c_{2}}(1+k_{s}^{-1})<c_{1}\tau/2-c_{3}$
$\implies t<\frac{k_{r}c_{2}\tau-2k_{r}c_{2}c_{3}/c_{1}}{2(1+k_{s}^{-1})}.$
Assuming that this bound on $t$ holds, then
$\exp\left(\epsilon\frac{(c_{1}/2+2c_{2})(c_{3}+s)}{4\eta c_{1}c_{2}}-\frac{\epsilon t}{8\eta c_{2}}+\frac{\epsilon r}{2c_{1}\eta}+\log\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}\right)\\ \leq\exp\left(\left[\frac{5+4k_{s}-k_{r}k_{s}}{8k_{r}k_{s}}\right]\frac{t\epsilon}{c_{2}\eta}+\log\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}+\frac{5c_{3}}{8c_{1}}\right).$ (B.14)
We next simplify the logarithmic term. First suppose that $r=tc_{1}/(k_{r}c_{2})$, then,
$\log\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}$ $=\log\frac{(2c_{2}\tau+tc_{1}/k_{r}c_{2}+c_{3})^{d}-t^{d}}{(tc_{1}/k_{r}c_{2})^{d}}$
$=d\log\frac{k_{r}c_{2}}{c_{1}}+\log\frac{(2c_{2}\tau+tc_{1}/k_{r}c_{2}+c_{3})^{d}-t^{d}}{t^{d}}$
$\leq d\log\frac{k_{r}c_{2}}{c_{1}}+d\log\frac{2c_{2}\tau+tc_{1}/k_{r}c_{2}+c_{3}}{t}$
$=d\log\frac{k_{r}c_{2}}{c_{1}}+d\log\left(\frac{2c_{2}\tau+c_{3}}{t}+\frac{c_{1}}{k_{r}c_{2}}\right)$
$=d\log\left(\frac{2k_{r}c_{2}^{2}\tau+c_{3}k_{r}c_{2}}{tc_{1}}+1\right).$ (B.15)
Now, from Lemma B.2, if $t\geq 2c_{2}\tau+s+c_{3}$, then $\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t\big{|}\tilde{\theta}\neq\perp\right)=0$ and so we consider $t<2c_{2}\tau+s+c_{3}$. Based on the assumed values for $r$ and $s$ this gives
$t<2c_{2}\tau+\frac{tc_{1}}{k_{s}k_{r}c_{2}}+c_{3}\implies t<\frac{2k_{s}k_{r}\tau c_{2}^{2}+c_{3}k_{s}k_{r}c_{2}}{k_{r}k_{s}c_{2}-c_{1}},$
whenever $k_{r}k_{s}c_{2}-c_{1}>0$. This holds trivially for any $k_{r}k_{s}\geq 1$. In order to ensure that the bounds discussed holds, we require that this upper bound on $t$ is less than the previous upper bound on $t$ to account for all values of $t$. Thus, using the fact that $k_{r}k_{s}\geq 1$, we have that
$\frac{2k_{s}k_{r}\tau c_{2}^{2}+c_{3}k_{s}k_{r}c_{2}}{k_{r}k_{s}c_{2}-c_{1}}\leq\frac{k_{r}c_{2}\tau-2k_{r}c_{2}c_{3}/c_{1}}{2(1+k_{s}^{-1})},$
whenever,
$\frac{2k_{s}k_{r}\tau c_{2}^{2}+c_{3}k_{s}k_{r}c_{2}}{k_{r}k_{s}c_{2}-c_{1}}\leq\frac{k_{r}c_{2}\tau-2k_{r}c_{2}c_{3}/c_{1}}{2(1+k_{s}^{-1})}$
$\implies\frac{2k_{s}\tau c_{2}+c_{3}k_{s}}{k_{r}k_{s}c_{2}-c_{1}}\leq\frac{\tau-2c_{3}/c_{1}}{2(1+k_{s}^{-1})}$
$\implies\frac{2(k_{s}+1)\tau c_{2}+c_{3}(k_{s}+1)}{k_{r}k_{s}c_{2}-c_{1}}+\frac{c_{3}}{c_{1}}\leq\tau/2$
$\implies\frac{c_{3}(k_{s}+1)}{k_{r}k_{s}c_{2}-c_{1}}+\frac{c_{3}}{c_{1}}\leq\tau\frac{k_{r}k_{s}c_{2}-c_{1}-4(k_{s}+1)c_{2}}{2(k_{r}k_{s}c_{2}-c_{1})}$
$\implies\tau\geq\frac{2(k_{r}k_{s}c_{2}-c_{1})}{k_{r}k_{s}c_{2}-c_{1}-4(k_{s}+1)c_{2}}\left(\frac{c_{3}(k_{s}+1)}{k_{r}k_{s}c_{2}-c_{1}}+\frac{c_{3}}{c_{1}}\right).$
Now, say we take $k_{s}=1$ and $k_{r}=10$, then
$\tau\geq\frac{20-2(c_{1}/c_{2})}{2-(c_{1}/c_{2})}\left(\frac{2c_{3}}{10c_{2}-c_{1}}+\frac{c_{3}}{c_{1}}\right),$
for which we can simplify this to the requirement that
$\tau\geq 20\left(\frac{2c_{3}}{10c_{2}-c_{1}}+\frac{c_{3}}{c_{1}}\right),$
Thus, for this choice of $r$ and $s$, (B.14) becomes
$\exp\left(\frac{\epsilon(c_{1}/2+2c_{2})(c_{3}+s)}{4\eta c_{1}c_{2}}-\frac{\epsilon t}{8\eta c_{2}}+\frac{\epsilon r}{2c_{1}\eta}+\log\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}+\frac{5c_{3}}{8c_{1}}\right)\\ \leq\exp\left(\left[\frac{5+4k_{s}-k_{r}k_{s}}{8k_{r}k_{s}}\right]\frac{t\epsilon}{c_{2}\eta}+d\log\left(\frac{2k_{r}c_{2}^{2}\tau}{tc_{1}}+1\right)+\frac{5c_{3}}{8c_{1}}\right).$
As a result, the above argumentation gives consideration of
$t<\frac{k_{r}c_{2}\tau-2k_{r}c_{2}c_{3}/c_{1}}{2(1+k_{s}^{-1})}.$ (B.16)
Using the derived upper bound, for $t>1/n$,
$\exp\left(\left[\frac{5+4k_{s}-k_{r}k_{s}}{8k_{r}k_{s}}\right]\frac{t\epsilon}{c_{2}\eta}+d\log\left(\frac{2k_{r}c_{2}^{2}\tau}{tc_{1}}+1\right)+\frac{5c_{3}}{8c_{1}}\right)\\ \leq\exp\left(\left[\frac{5+4k_{s}-k_{r}k_{s}}{8k_{r}k_{s}}\right]\frac{t\epsilon}{c_{2}\eta}+d\log\left(\frac{2nk_{r}c_{2}^{2}\tau}{c_{1}}+1\right)+\frac{5c_{3}}{8c_{1}}\right).$
Bounding this above by $\gamma/3$ yields
$\exp\left(\left[\frac{5+4k_{s}-k_{r}k_{s}}{8k_{r}k_{s}}\right]\frac{t\epsilon}{c_{2}\eta}+d\log\left(\frac{2nk_{r}c_{2}^{2}\tau}{c_{1}}+1\right)\right)\leq\frac{\gamma}{3}\exp\left(-\frac{5c_{3}}{8c_{1}}\right)$
$\implies\left[\frac{k_{r}k_{s}-5-4k_{s}}{8k_{r}k_{s}}\right]\frac{t\epsilon}{c_{2}\eta}\geq-\log\left(\frac{\gamma}{3}\exp\left(-\frac{5c_{3}}{8c_{1}}\right)\left(\frac{2nk_{r}c_{2}^{2}\tau}{c_{1}}+1\right)^{-d}\right)$
$\implies t>\frac{\eta c_{2}}{\epsilon}\left[\frac{8k_{r}k_{s}}{k_{r}k_{s}-5-4k_{s}}\right]\log\left(\frac{3}{\gamma}\exp\left(\frac{5c_{3}}{8c_{1}}\right)\left(\frac{2nk_{r}c_{2}^{2}\tau}{c_{1}}+1\right)^{d}\right)$
$\implies t>80\frac{\eta c_{2}}{\epsilon}\log\left(\frac{3}{\gamma}\exp\left(\frac{5c_{3}}{8c_{1}}\right)\left(\frac{20nc_{2}^{2}\tau}{c_{1}}+1\right)^{d}\right)$
Therefore, this reduces to
$t\gtrsim\frac{\eta c_{2}}{\epsilon}\left[\log(1/\gamma)\vee d\log\left(\frac{20c_{2}^{2}n\tau}{c_{1}}+1\right)\vee c_{3}/c_{1}\right]\vee 1/n.$
Next we bound $\zeta_{n,d}(s\wedge c_{3}\wedge c_{1}/2)\leq\gamma/3$. By assumption, $\gamma>3\zeta_{n,d}(c_{1}/2\wedge c_{3})$, and so we consider only $\zeta_{n,d}(s)$ for $s<c_{1}/2\wedge c_{3}$. To this end, using the fact that $\zeta_{n,d}$ are non-increasing, we have that
$\zeta_{n,d}(s)\leq\gamma/3\implies t\gtrsim\frac{c_{2}}{c_{1}}\zeta_{n,d}^{-1}(\gamma/3).$
This completes the proof.
Lemma B.2.
If Conditions 2 and 3 hold, then for all $\eta,\epsilon,\delta>0$, $t>0$, and $n,d\geq 1$ then
$\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t\big{|}\tilde{\theta}\neq\perp\right)\\ \lesssim\inf_{\begin{subarray}{c}0<s<c_{1}\tau/2-c_{3}\\ 0<r<c_{1}\tau/2-s-c_{3}\end{subarray}}\Bigg{[}\exp\left(\epsilon\frac{(c_{1}/2+2c_{2})(c_{3}+s)}{4\eta c_{1}c_{2}}-\frac{\epsilon t}{8\eta c_{2}}+\frac{\epsilon r}{2c_{1}\eta}+\log\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}\right)\\ +\zeta_{n,d}\left(s\wedge c_{3}\wedge c_{1}/2\right)\Bigg{]}.$
Proof of Lemma B.2.
Let $\beta=\epsilon/4\eta$ and define the events $E_{6,s}=\{S_{n}(\mu)\leq s\}$ and
$E_{7}=\left\{\inf_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})\geq c_{1}/2\right\}\cap\left\{\sup_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})\leq 2c_{2}\right\}\cap\left\{\hat{\alpha}_{\mu}\leq 4c_{3}\right\}.$
Condition 2 directly gives that $\Pr\left(E_{6,s}^{c}\right)\lesssim\zeta_{n,d}(s)$. Using this, the identities established in (B.2), (B.3), and (B.4), and noting that $c_{1}\leq c_{2}$, yields that for all $s\geq 0$, we have
$\Pr\left((E_{7}\cap E_{6,s})^{c}\right)\lesssim\zeta_{n,d}(s\wedge c_{3}\wedge c_{1}/2).$ (B.17)
Next, condition on $E_{7}\cap E_{6,s}$. For any $\lVert x-\theta\rVert\geq t$,
$O(x,\hat{\nu})$ $=\sup_{u\in\mathbb{S}^{d-1}}\frac{x^{\top}u-\theta^{\top}u+\theta^{\top}u-\hat{\mu}_{u}}{\sigma(\hat{\nu}_{u})}$
$\geq\frac{\sup_{u\in\mathbb{S}^{d-1}}(x-\theta)^{\top}u-\sup_{u\in\mathbb{S}^{d-1}}|\hat{\mu}_{u}-\theta^{\top}u|}{\sup_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}}$
$=\frac{\lVert x-\theta\rVert-\sup_{u\in\mathbb{S}^{d-1}}|\hat{\mu}_{u}-\mu_{u}+\mu_{u}-\theta^{\top}u|}{\sup_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}}$
$\geq\frac{\lVert x-\theta\rVert-\sup_{u\in\mathbb{S}^{d-1}}|\hat{\mu}_{u}-\mu_{u}|-\sup_{u\in\mathbb{S}^{d-1}}|\mu_{u}-\theta^{\top}u|}{\sup_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}}$
$\geq\frac{t-c_{3}-s}{2c_{2}}.$ (B.18)
The last inequality relies on $E_{6,s}$ and Condition 3. A similar argument shows that, for any $r>0$ and $x\in\mathbb{R}^{d}$ such that $\lVert x-\theta\rVert\leq r$,
$O(x,\hat{\nu})$ $=\sup_{u\in\mathbb{S}^{d-1}}\frac{x^{\top}u-\theta^{\top}u+\theta^{\top}u-\hat{\mu}_{u}}{\sigma(\hat{\nu}_{u})}$
$\leq\frac{\sup_{u\in\mathbb{S}^{d-1}}(x^{\top}u-\theta^{\top}u)+\sup_{u\in\mathbb{S}^{d-1}}(\theta^{\top}u-\hat{\mu}_{u})}{\inf_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})}$
$\leq\frac{r+\sup_{u\in\mathbb{S}^{d-1}}|\theta^{\top}u-\mu_{u}|+\sup_{u\in\mathbb{S}^{d-1}}|\mu_{u}-\hat{\mu}_{u}|}{c_{1}/2}$
$\leq\frac{r+c_{3}+s}{c_{1}/2}.$ (B.19)
Therefore, taking $B_{r}(x)$ to be the ball of radius $r$ centered on $x$, then
$B_{c_{1}\tau/2-s-c_{3}}(\theta)\subset A_{\tau,\hat{\nu}}=\{x\mid O(x,\hat{\nu})<\tau\}\subset B_{2c_{2}\tau+s+c_{3}}(\theta).$ (B.20)
Here, the first inclusion follows from (B.19) and the second from (B.18).
Next, note that by definition
[…]
Moreover, if $t\geq 2c_{2}\tau+s+c_{3}$, (B.20) indicates that $B_{t}^{c}(\theta)\cap A_{\tau,\hat{\nu}}=\emptyset$ and so
[…]
As such, we restrict $t<2c_{2}\tau+s+c_{3}$. Under this restriction, applying (B.18) we find that
[…]
$=\exp{\left(\frac{-\beta(t-c_{3}-s)}{2c_{2}}\right)}\operatorname{vol}(\{t\leq\lVert x\rVert\leq 2c_{2}\tau+s+c_{3}\}).$
Similarly, an application of (B.19) gives
[…]
$\geq\sup_{0<r\leq c_{1}\tau/2-s-c_{3}}\exp{\left(\frac{-\beta(r+c_{3}+s)}{c_{1}/2}\right)}\operatorname{vol}(B_{r}(\theta)).$
Combining these two bounds gives
$\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t|E_{7}\cap E_{6,s}\right)$
$\leq\frac{\exp{\left(\frac{-\beta(t-c_{3}-s)}{2c_{2}}\right)}\operatorname{vol}(\{t\leq\lVert x\rVert\leq 2c_{2}\tau+s+c_{3}\})}{\sup_{0<r\leq c_{1}\tau/2-s-c_{3}}\exp{\left(\frac{-\beta(r+c_{3}+s)}{c_{1}/2}\right)}\operatorname{vol}(B_{r}(\theta))}$
$=\inf_{0<r\leq c_{1}\tau/2-s-c_{3}}\frac{\exp{\left(\frac{-\beta(t-c_{3}-s)}{2c_{2}}\right)}\operatorname{vol}(B_{t}^{c}(0)\cap B_{2c_{2}\tau+s+c_{3}}(0))}{\exp{\left(\frac{-\beta(r+c_{3}+s)}{c_{1}/2}\right)}\operatorname{vol}(B_{r}(\theta))}$
$\lesssim\inf_{0<r\leq c_{1}\tau/2-s-c_{3}}\exp{\left(-\beta\frac{c_{1}/2(t-c_{3}-s)-2c_{2}(r+c_{3}+s)}{c_{2}c_{1}}\right)}\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}$
$=\inf_{0<r\leq c_{1}\tau/2-s-c_{3}}\exp{\left(\beta\left(\frac{(c_{1}/2+2c_{2})(c_{3}+s)}{c_{1}c_{2}}-\frac{t}{2c_{2}}+\frac{2r}{c_{1}}\right)\right)}\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}.$
This bound holds for all $0<s<c_{1}\tau/2-c_{3}$, and so we can take the least upper bound on the range. Combining this and re-writing the expression gives
$\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t|E_{7}\cap E_{6,s}\right)\lesssim\inf_{\begin{subarray}{c}0<r\leq c_{1}\tau/2-s-c_{3}\\ 0<s<c_{1}\tau/2-c_{3}\end{subarray}}\exp\left(\beta\left(\frac{(c_{1}/2+2c_{2})(c_{3}+s)}{c_{1}c_{2}}-\frac{t}{2c_{2}}+\frac{2r}{c_{1}}\right)\right.\\ \left.+\log\left(\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}\right)\right).$
Since $\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t\right)\leq\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t|E_{7}\cap E_{6,s}\right)+P((E_{7}\cap E_{6,s})^{c})$, combining this result with (B.17) gives the desired result. ∎
Lemma B.3.
For all $\tau>0$, $n,d\geq 1$, $\hat{\nu}\in\widehat{\mathcal{M}}(n,d)$ and $k\in[n]$
$\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k)\\ x\in A_{\tau,\tilde{\nu}}\cup A_{\tau,\hat{\nu}}\end{subarray}}|O(x,\hat{\nu})-O(x,\tilde{\nu})|\leq\frac{\tau\tilde{S}_{n,k}(\sigma)+\tilde{S}_{n,k}(\mu)}{\inf_{u\in\mathbb{S}^{d-1}}b_{k,u}^{-}}\coloneqq\Delta_{k}.$ (B.21)
Proof.
We use the fact that
$\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k)\\ x\in A_{\tau,\tilde{\nu}}\cup A_{\tau,\hat{\nu}}\end{subarray}}|O(x,\hat{\nu})-O(x,\tilde{\nu})|\leq\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k)\\ u\in\mathbb{S}^{d-1}\\ x\in A_{\tau,\tilde{\nu}}\cup A_{\tau,\hat{\nu}}\end{subarray}}|O_{u}(x,\hat{\nu})-O_{u}(x,\tilde{\nu})|,$
and bound the right-hand side. Observe that for all $\tau>0$ and all $\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k)$ it holds that
$\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ x\in A_{\tau,\hat{\nu}}\end{subarray}}|O_{u}(x,\hat{\nu})-O_{u}(x,\tilde{\nu})|$ $=\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ x\in A_{\tau,\hat{\nu}}\end{subarray}}\left|\frac{x^{\top}u-\hat{\mu}_{u}}{\hat{\sigma}_{u}}-\frac{x^{\top}u-\mu(\tilde{\nu}_{u})}{\sigma(\tilde{\nu}_{u})}\right|$
$\leq\sup_{u\in\mathbb{S}^{d-1}}\frac{1}{\hat{\sigma}_{u}\sigma(\tilde{\nu}_{u})}\left(\left(\sup_{x\in A_{\tau,\tilde{\nu}}}|x^{\top}u-\hat{\mu}_{u}|\right)\tilde{S}_{n,k}(\sigma)+\sigma(\hat{\nu}_{u})\tilde{S}_{n,k}(\mu)\right)$
$\leq\frac{\tau\tilde{S}_{n,k}(\sigma)+\tilde{S}_{n,k}(\mu)}{\inf_{u\in\mathbb{S}^{d-1}}b^{-}_{k,u}}.$
Using the same logic, for all $\tau>0$ and all $\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k)$ we can write:
$\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ x\in A_{\tau,\tilde{\nu}}\end{subarray}}|O_{u}(x,\hat{\nu})-O_{u}(x,\tilde{\nu})|$ $\leq\frac{\tau\tilde{S}_{n,k}(\sigma)+\tilde{S}_{n,k}(\mu)}{\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}}\leq\frac{\tau\tilde{S}_{n,k}(\sigma)+\tilde{S}_{n,k}(\mu)}{\inf_{u\in\mathbb{S}^{d-1}}b^{-}_{k,u}}.$
Combining these two results yields that, for all $\tau>0$ and all $\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k)$, it holds that
$\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ x\in A_{\tau,\tilde{\nu}}\cup A_{\tau,\hat{\nu}}\end{subarray}}|O_{u}(x,\hat{\nu})-O_{u}(x,\tilde{\nu})|\leq\frac{\tau\tilde{S}_{n,k}(\sigma)+\tilde{S}_{n,k}(\mu)}{\inf_{u\in\mathbb{S}^{d-1}}b_{k,u}^{-}}.\qed$
Let $\hat{\alpha}_{\mu}=\sup_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}-\inf_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}$.
Lemma B.4.
For $\phi=O$, and all $\hat{\nu}\in\widehat{\mathcal{M}}_{1}(n,d)$ and $\tau,\eta,y>0$ such that $\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-y-2\eta)-\hat{\alpha}_{\mu}>0$, it holds that
$\frac{\operatorname{vol}(A_{\tau+2\eta,\hat{\nu}}\cap A^{c}_{\tau-2\eta,\hat{\nu}})}{\operatorname{vol}(A_{\tau-y-2\eta,\hat{\nu}})}\leq\left(\frac{\sup_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau+2\eta)+\hat{\alpha}_{\mu}}{\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-y-2\eta)-\hat{\alpha}_{\mu}}\right)^{d}-\left(\frac{\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-2\eta)-\hat{\alpha}_{\mu}}{\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-y-2\eta)-\hat{\alpha}_{\mu}}\right)^{d}.$
Proof.
Let $a(t)=\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}t-\hat{\alpha}_{\mu}$ and $b(t)=\sup_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}t+\hat{\alpha}_{\mu}$. We first show that $\operatorname{vol}(B_{a(t)})\leq\operatorname{vol}(A_{t,\hat{\nu}})\leq\operatorname{vol}(B_{b(t)}).$ Beginning with the upper bound, for any $z\in\mathbb{R}^{d}$,
$A_{t,\hat{\nu}}$ =$=\{x\colon\sup_{u\in\mathbb{S}^{d-1}}\frac{|x^{\top}u-\hat{\mu}_{u}|}{\hat{\sigma}_{u}}\leq t\}\subset\{x\colon\sup_{u\in\mathbb{S}^{d-1}}|x^{\top}u-z^{\top}u|\leq\sup_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}t+\sup_{u\in\mathbb{S}^{d-1}}|z^{\top}u-\hat{\mu}_{u}|\}.$
In particular, taking $z=v\hat{\mu}_{v}$ for $v\in\mathbb{S}^{d-1}$,
$A_{t,\hat{\nu}}$$\subset\{|x^{\top}u-\hat{\mu}_{v}v^{\top}u|\leq b(t)\}=\{x\colon\lVert x-\hat{\mu}_{v}v\rVert\leq b(t)\}=B_{b(t)}(\hat{\mu}_{v}v).$
It follows that $\operatorname{vol}(A_{t,\hat{\nu}})\leq\operatorname{vol}(B_{b(t)}).$ Next, we show the analogous lower bound. To this end, note that for any $z\in\mathbb{R}^{d}$
$A_{t,\hat{\nu}}$ =$=\{x\colon\sup_{u\in\mathbb{S}^{d-1}}\frac{|x^{\top}u-\hat{\mu}_{u}|}{\hat{\sigma}_{u}}\leq t\}\supset\{x\colon\sup_{u\in\mathbb{S}^{d-1}}|x^{\top}u-z^{\top}u|\leq\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}t-\sup_{u\in\mathbb{S}^{d-1}}|z^{\top}u-\hat{\mu}_{u}|\}.$
Take $z=v\hat{\mu}_{v}$, where $v\in\mathbb{S}^{d-1}$, yielding $A_{t,\hat{\nu}}\supset\{x\colon\lVert x-v\hat{\mu}_{v}\rVert\leq a(t)\}=B_{a(t)}(v\hat{\mu}_{v}).$ It follows that $\operatorname{vol}(A_{t,\hat{\nu}})\geq\operatorname{vol}(B_{a(t)}).$ Now, we apply the upper and lower bounds, in conjunction with the assumption that $a(\tau-y-2\eta)=\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-y-2\eta)-\hat{\alpha}_{\mu}>0$, which results in
$\frac{\operatorname{vol}(A_{\tau+2\eta,\hat{\nu}}\cap A^{c}_{\tau-2\eta,\hat{\nu}})}{\operatorname{vol}(A_{\tau-y-2\eta,\hat{\nu}})}$ $\leq\frac{\operatorname{vol}(B_{b(\tau+2\eta)}\cap B^{c}_{a(\tau-2\eta)})}{\operatorname{vol}(B_{a(\tau-y-2\eta)})}$
$\leq\left(\frac{\sup_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau+2\eta)+\hat{\alpha}_{\mu}}{\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-y-2\eta)-\hat{\alpha}_{\mu}}\right)^{d}$
$\hskip 122.85876pt-\left(\frac{\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-2\eta)-\hat{\alpha}_{\mu}}{\inf_{u\in\mathbb{S}^{d-1}}\hat{\sigma}_{u}(\tau-y-2\eta)-\hat{\alpha}_{\mu}}\right)^{d}.\qed$
Appendix C Proofs from Section 3.2
Let $\lambda_{t,q_{1},q_{2}}=\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ z\in(\xi_{q_{1},u}-t,\xi_{q_{2},u}+t)\end{subarray}}f_{\nu_{u}}(z)$ and $\hat{\xi}_{q,u}=F^{-1}_{\hat{\nu}_{u}}(q)$.
Lemma C.1.
If $\nu$ is absolutely continuous, then for all for $10<q_{1}\leq q_{2}<$ 1, $n,d\geq 1$ and $t>0$ it holds that
$\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}\sup_{x\in[q_{1},q_{2}]}|\xi_{x,u}-\hat{\xi}_{x,u}|\geq t\right)\leq\left(\frac{cn}{d}\right)^{d}e^{-2nt^{2}\lambda_{t,q_{1},q_{2}}^{2}}.$
Proof.
Note that
$\Pr\Big{(}\sup_{\begin{subarray}{c}x\in[q_{1},q_{2}]\\ u\in\mathbb{S}^{d-1}\end{subarray}}|\xi_{x,u}-\hat{\xi}_{x,u}|\geq t\Big{)}=\Pr\Big{(}\bigcup_{\begin{subarray}{c}x\in[q_{1},q_{2}]\\ u\in\mathbb{S}^{d-1}\end{subarray}}\left(\{\hat{\xi}_{x,u}-\xi_{x,u}\geq t\}\cup\{\xi_{x,u}-\hat{\xi}_{x,u}\geq t\}\right)\Big{)}.$ (C.1)
Furthermore, a direct consequence of the mean value theorem is that
$\{\hat{\xi}_{x,u}-\xi_{x,u}\geq t\}$ =$=\{x\geq F_{\hat{\nu}_{u}}(\xi_{x,u}+t)\}$
=$=\{F_{\nu_{u}}(\xi_{x,u}+t)-F_{\hat{\nu}_{u}}(\xi_{x,u}+t)\geq F_{\nu_{u}}(\xi_{x,u}+t)-x\}$
$\subset\{\sup_{\begin{subarray}{c}y\in[0,1]\\ u\in\mathbb{S}^{d-1}\end{subarray}}|F_{\nu_{u}}(y)-F_{\hat{\nu}_{u}}(y)|\geq t\inf_{\begin{subarray}{c}z\in[q_{1},q_{2}]\\ u\in\mathbb{S}^{d-1}\end{subarray}}f_{\nu_{u}}(\xi_{z,u}+t)\}.$
A similar argument yields
$\{\xi_{x,u}-\hat{\xi}_{x,u}\geq t\}$$\subset\{\sup_{\begin{subarray}{c}y\in[0,1]\\ u\in\mathbb{S}^{d-1}\end{subarray}}|F_{\nu_{u}}(y)-F_{\hat{\nu}_{u}}(y)|\geq t\inf_{\begin{subarray}{c}z\in[q_{1},q_{2}]\\ u\in\mathbb{S}^{d-1}\end{subarray}}f_{\nu_{u}}(\xi_{z,u}-t)\}.$
Combining these two set inequalities with (C.1) and applying Talagrand’s inequality [23] results in
$\Pr\Big{(}\sup_{\begin{subarray}{c}x\in[q_{1},q_{2}]\\ u\in\mathbb{S}^{d-1}\end{subarray}}|\xi_{x,u}-\hat{\xi}_{x,u}|\geq t\Big{)}$ $\leq\Pr\Big{(}\sup_{\begin{subarray}{c}y\in[0,1]\\ u\in\mathbb{S}^{d-1}\end{subarray}}|F_{\nu_{u}}(y)-F_{\hat{\nu}_{u}}(y)|\geq t\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ z\in(\xi_{q_{1},u}-t,\xi_{q_{2},u}+t)\end{subarray}}f_{\nu_{u}}(z)\Big{)}$
$=\Pr\Big{(}\sup_{\begin{subarray}{c}y\in[0,1]\\ u\in\mathbb{S}^{d-1}\end{subarray}}|F_{\nu_{u}}(y)-F_{\hat{\nu}_{u}}(y)|\geq t\lambda_{t,q_{1},q_{2}}\Big{)}$
$\leq\left(\frac{cn}{d}\right)^{d}e^{-2nt^{2}\lambda_{t,q_{1},q_{2}}^{2}}.\qed$
For $\mathbb{X}_{n}\in\mathbf{D}_{n\times d}$ and $u\in\mathbb{S}^{d-1}$, let $(X^{\top}u)_{(k)}$ denote the $k$th order statistic of the sample $X_{1}^{\top}u,\ldots,X_{n}^{\top}u$.
Lemma C.2.
If, for a given $20<q\leq 1/$ 2, Condition 4 holds for $(q,\nu,k)$, then there exists a universal constant $c>0$ such that for all $d,n\geq 1$ and all $C>0$, it holds that
$\Pr\Big{(}\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ p\in[q,1-q-k/n]\end{subarray}}(X^{\top}u)_{(\lfloor pn\rfloor+k)}-(X^{\top}u)_{(\lfloor pn\rfloor)}\geq C/n\Big{)}\lesssim\exp\left(-nr^{2}M^{2}+d\log(cn/d)\right)\\ +\exp\left(-\frac{3MC}{4k}+d\log 9+\log n(1-2q+1/n)\right).$
Proof.
Let $m=\lfloor qn\rfloor$. Now, define
$E_{8}=\bigcap_{u\in\mathbb{S}^{d-1}}\left\{\xi_{q,u}-r\leq(X^{\top}u)_{(m)}\leq(X^{\top}u)_{(m+k)}\leq\xi_{q,u}+r\right\}.$
Now, noting that, for a given $j\in[n]$, $(X^{\top}u)_{(j)}=\hat{\xi}_{j/n,u}$, applying Lemma C.1 with $q_{1}=q$, $q_{2}=q+k/n$ and $t=r$ in conjunction with Condition 4 yields
$\Pr\left(E_{8}^{c}\right)\lesssim\exp\left(-nr^{2}\lambda_{r,q,q+k/n}^{2}+d\log(cn/d)\right)\leq\exp\left(-nr^{2}M^{2}+d\log(cn/d)\right).$ (C.2)
Now, let
$E_{9}=\left\{\max_{m\leq j<n-m}\sup_{u\in\mathbb{S}^{d-1}}\left[(X^{\top}u)_{(j+1)}-(X^{\top}u)_{(j)}\right]\geq\frac{C}{kn}\right\},$
and note that
$\Pr\Big{(}\ \sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ p\in[q,1-q-k/n]\end{subarray}}(X^{\top}u)_{(\lfloor pn\rfloor+k)}-(X^{\top}u)_{(\lfloor pn\rfloor)}\geq\frac{C}{n}\Big{)}\leq\Pr\left(E_{9}|E_{8}\right)+\Pr\left(E_{8}^{c}\right).$ (C.3)
We now consider the event $E_{9}$. Recall that the set $\mathcal{N}_{\epsilon}$ is an $\epsilon$-net on $\mathbb{S}^{d-1}$ if for all $u\in\mathbb{S}^{d-1}$, there exists $v\in\mathcal{N}_{\epsilon}$ such that $\lVert u-v\rVert\leq\epsilon$. Letting $\mathcal{N}_{\epsilon}$ be an arbitrary $\epsilon$-net on $\mathbb{S}^{d-1}$, it is straightforward (see, e.g., [25, Lemma 4.4.1]) to show that for any $i,j\in[n]$,
$\sup_{u\in\mathbb{S}^{d-1}}|X_{i}^{\top}u-X_{j}^{\top}u|\leq\frac{1}{1-\epsilon}\sup_{u\in\mathcal{N}_{\epsilon}}|X_{i}^{\top}u-X_{j}^{\top}u|.$
Furthermore, there exists $\mathcal{N}_{1/4}$ such that $|\mathcal{N}_{1/4}|\leq 9^{d}$ [25, see Corollary 4.2.13]. Applying these two facts, in conjunction with two applications of subadditivity of probability measures, yields
$\Pr\left(E_{8}\right)$ $\leq\Pr\left(\max_{m\leq j<n-m}\sup_{u\in\mathcal{N}_{1/4}}(X^{\top}u)_{(j+1)}-(X^{\top}u)_{(j)}\geq\frac{3C}{4n}\right)$
$\leq 9^{d}\sup_{u\in\mathcal{N}_{1/4}}\Pr\left(\max_{m\leq j<n-m}(X^{\top}u)_{(j+1)}-(X^{\top}u)_{(j)}\geq\frac{3C}{4n}\right)$
$\leq 9^{d}(n-2m+1)\sup_{\begin{subarray}{c}u\in\mathcal{N}_{1/4}\\ m\leq j<n-m\end{subarray}}\Pr\left((X^{\top}u)_{(j+1)}-(X^{\top}u)_{(j)}\geq\frac{3C}{4n}\right)$ (C.6)
$\coloneqq e^{d\log 9+\log(n-2m+1)}\sup_{\begin{subarray}{c}u\in\mathcal{N}_{1/4}\\ m\leq j<n-m\end{subarray}}\Pr\left(A_{3,u,j}\right).$ (C.9)
Now, observe that for any $m\leq j<n-m$ and $u\in\mathbb{S}^{d-1}$, using Condition 4, together with the mean value theorem, imply that on $E_{8}$ and $A_{3,u,j}$ it holds that
$\frac{F_{\nu_{u}}((X^{\top}u)_{(j+1)})-F_{\nu_{u}}((X^{\top}u)_{(j)})}{(X^{\top}u)_{(j+1)}-(X^{\top}u)_{(j)}}\geq M\implies F_{\nu_{u}}((X^{\top}u)_{(j+1)})-F_{\nu_{u}}((X^{\top}u)_{(j)})\geq\frac{3C}{4n}M.$ (C.10)
Let $U_{(j),u}=F_{\nu_{u}}((X^{\top}u)_{(j)})$. Now, recall that for any set of $n$ independent standard uniform random variables $V_{1},\ldots,V_{n}$, for any $j\in[n-1]$, it holds that $V_{(j+1)}-V_{(j)}\stackrel{{\scriptstyle\text{d}}}{{=}}V_{(1)}$ and $V_{(1)}\sim Beta(1,n)$ [see, e.g., 19]. These two facts, combined inequality (C.10), results in
$\sup_{\begin{subarray}{c}u\in\mathcal{N}_{1/4}\\ m\leq j<n-m\end{subarray}}\Pr\left(A_{3,u}|E_{8}\right)$ $\leq\sup_{\begin{subarray}{c}u\in\mathcal{N}_{1/4}\\ m\leq j<n-m\end{subarray}}\Pr\left(U_{(j+1),u}-U_{(j),u}\geq 3CM/4n\right)$ (C.15)
$\leq\sup_{u\in\mathcal{N}_{1/4}}\Pr\left(U_{(1),u}\geq 3CM/4n\right)$
$\leq(1-(3CM/4n)\wedge 1)^{n}$
$\leq e^{-3MC/4k}.$ (C.16)
Combining (C.2)–(C.16) yields
$\Pr\Big{(}\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ p\in[q,1-q-k]\end{subarray}}(X^{\top}u)_{(\lfloor pn\rfloor+k)}-(X^{\top}u)_{(\lfloor pn\rfloor)}\geq\frac{C}{n}\Big{)}\lesssim\exp\left(-nr^{2}M^{2}+d\log(cn/d)\right)\\ \exp\left(-3MC/4k+d\log 9+\log(n-2m+1)\right).\qed$
Lemma C.3.
Let $\mu$ be the median functional. If Condition 4 holds for $(1/2,\nu,k)$, then there exists a universal constant $c>0$ such that for all $d,n\geq 1$ and $t>0$ it holds that
$\Pr\left(\tilde{S}_{n,k}(\mu)\geq t\right)\lesssim\exp\left(-nr^{2}M^{2}+d\log(cn/d)\right)+\exp\left(-3MC/4k+d\log 9+\log n\right).$
Proof of Lemma C.3.
Observe that
$\tilde{S}_{n,k}(\mu)\leq\sup_{u\in\mathbb{S}^{d-1}}|(X^{\top}u)_{(\lfloor(n+1)/2\rfloor+k)}-(X^{\top}u)_{(\lfloor(n+1)/2\rfloor)}|.$ (C.17)
Therefore, applying Lemma C.2 with $q=1/2$ and $k=k$ gives the result.
Lemma C.4.
Suppose that Condition 4 holds for $(q^{\prime},\nu,k)$ with $q^{\prime}<1/4$ and that there exists $q^{\prime}<q<1/4$ such that $k<n(q-q^{\prime})/2$. Then for all $d,n\geq 1$ and $C>0$ it holds that
$\Pr\Big{(}\tilde{S}_{n,k}(\sigma)\geq C/n\Big{)}\lesssim\exp\left(-nr^{2}M^{2}+d\log(cn/d)\right)\\ +\exp\left(-3MC/8k+d\log 9+\log(n(1-2q^{\prime}+1/n))\right).$
Proof.
Let
$E_{10}=\Big{\{}\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ p\in[q^{\prime},1-q^{\prime}-k/n]\end{subarray}}(X^{\top}u)_{(\lfloor pn\rfloor+k)}-(X^{\top}u)_{(\lfloor pn\rfloor)}<C/2n\Big{\}}.$
A direct application of Lemma C.2 gives that for all $u\in\mathbb{S}^{d-1}$ and $\tilde{\nu}_{u}\in\mathcal{M}(k,\hat{\nu})$,
$\Pr\left(E_{10}^{c}\right)\leq\exp\left(-3MC/8k+d\log 9+\log(n(1-2q^{\prime}+1/n))\right)\\ +\exp\left(-nr^{2}M^{2}+d\log(cn/d)\right).$
It remains to show that on $E_{10}$, for all $u\in\mathbb{S}^{d-1}$ and $\tilde{\nu}_{u}\in\mathcal{M}(k,\hat{\nu})$, it holds that $\sigma(\hat{\nu}_{u})-C/n\leq\sigma(\tilde{\nu}_{u})\leq\sigma(\hat{\nu}_{u})+C/n$. We begin by proving the upper bound. To this end, (C.17) implies that on $E_{10}$ it holds that $\tilde{S}_{n,k}(\mu)<C/2n.$ This fact implies that
$[\hat{\mu}_{u}-\hat{\sigma}_{u},\hat{\mu}_{u}+\hat{\sigma}_{u}]\subset[\mu(\tilde{\nu}_{u})-\hat{\sigma}_{u}-C/2n,\mu(\tilde{\nu}_{u})+\hat{\sigma}_{u}+C/2n].$ (C.18)
Next, recall that, by definition, the median absolute deviation is equal to half the length of the smallest interval which is symmetric about the median and which contains $\lfloor(n+1)/2\rfloor$ of the observations. Therefore, $\lfloor(n+1)/2\rfloor$ of the projected observations are contained in $[\hat{\mu}_{u}-\hat{\sigma}_{u},\hat{\mu}_{u}+\hat{\sigma}_{u}]$. Consequentially, (C.18) implies that on $E_{10}$, for all $u\in\mathbb{S}^{d-1}$ and $\tilde{\nu}_{u}\in\mathcal{M}(k,\hat{\nu})$, $\lfloor(n+1)/2\rfloor$ of the projected observations are also contained in $[\mu(\tilde{\nu}_{u})-\hat{\sigma}_{u}-C/2n,\mu(\tilde{\nu}_{u})+\hat{\sigma}_{u}+C/2n]$. This directly implies that for all $u\in\mathbb{S}^{d-1}$ and $\tilde{\nu}_{u}\in\mathcal{M}(k,\hat{\nu})$, it holds that $\sigma(\tilde{\nu}_{u})\leq\sigma(\hat{\nu}_{u})+C/2n$.
It remains to show that on $E_{10}$, $\sigma(\hat{\nu}_{u})-C/n\leq\sigma(\tilde{\nu}_{u})$. First, by definition, for all $u\in\mathbb{S}^{d-1}$ and $\tilde{\nu}_{u}\in\mathcal{M}(k,\hat{\nu})$, $\lfloor(n+1)/2\rfloor-k$ of the projected observations are contained in $[\mu(\tilde{\nu}_{u})-\sigma(\tilde{\nu}_{u}),\mu(\tilde{\nu}_{u})+\sigma(\tilde{\nu}_{u})]$. Next, note that for all $u\in\mathbb{S}^{d-1}$ there exists $q<p_{u}<1-q$ such that $\hat{\sigma}_{u}=|\mu(\hat{\nu}_{u})-\hat{\xi}_{p_{u},u}|$. To see this, take $q^{\prime}<q<1/4$, and note that by definition,
$\hat{\sigma}_{u}\leq|\hat{\xi}_{q,u}-\hat{\xi}_{1/2,u}|\vee|\hat{\xi}_{1-q,u}-\hat{\xi}_{1/2,u}|.$
(This is because $[\hat{\xi}_{q,u},\hat{\xi}_{1-q,u}]$ always contains at least half of the observations by definition ($q<1/4$) and, therefore, $\mu_{u}\pm|\hat{\xi}_{q,u}-\hat{\xi}_{1/2,u}|\vee|\hat{\xi}_{1-q,u}-\hat{\xi}_{1/2,u}|$ contains at least half of the observations.) Therefore, there exists $q<p_{u}<1-q$ such that $\hat{\sigma}_{u}=|\hat{\xi}_{1/2,u}-\hat{\xi}_{p_{u},u}|$.
Now, by assumption, $2k<n(q-q^{\prime})$, and so, for all $u\in\mathbb{S}^{d-1}$, there exists $q^{\prime}<p_{u}<1-q^{\prime}$ such that $\sigma(\tilde{\nu}_{u})=|\mu(\tilde{\nu}_{u})-\hat{\xi}_{p_{u},u}|$. (This is because $\hat{\xi}_{q,u}$ is at least $2k$ observations away from $\hat{\xi}_{q^{\prime},u}$.) Therefore, on $E_{10}$, $\lfloor(n+1)/2\rfloor$ of the observations lie within $[\mu(\hat{\nu}_{u})-\sigma(\tilde{\nu}_{u})-C/n,\mu(\hat{\nu}_{u})+\sigma(\tilde{\nu}_{u})+C/n]$, which immediately implies that $\sigma(\hat{\nu}_{u})-C/n\leq\sigma(\tilde{\nu}_{u})$.
Lemma C.5.
Let $\sigma$ be the interquartile range. If Condition 4 holds for $(1/4,\nu,k)$, then for all $d,n\geq 1$ and $t>0$ it holds that
$\Pr\left(\tilde{S}_{n,k}(\sigma)\geq t\right)\lesssim\exp\left(-3MC/8k+d\log 9+\log(n(1/2+1/n))\right)\\ +\exp\left(-nr^{2}M^{2}+d\log(cn/d)\right).$
Proof of Lemma C.5.
Observe that
$\tilde{S}_{n,k}(\sigma)\leq 2\sup_{u\in\mathbb{S}^{d-1}}|(X^{\top}u)_{(\lfloor(n+1)/4\rfloor+k)}-(X^{\top}u)_{(\lfloor(n+1)/4\rfloor)}|.$
Therefore, applying Lemma C.2 with $q=1/4$ and $k=k$ yields the result.
Appendix D Proofs and related lemmas from Section 5
[…]
Lemma D.1.
For any $\hat{\nu}\in\widehat{\mathcal{M}}_{1}(n,d)$, $\tau,\epsilon>0$, $0<\eta\leq\tau$ and $\delta\leq e^{\epsilon/2}/2$ if
$\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},1)\\ x\in A_{\tau,\hat{\nu}}\cup A_{\tau,\tilde{\nu}}\end{subarray}}|\phi(x,\hat{\nu})-\phi(x,\tilde{\nu})|\leq\eta,$ (D.1)
and
$\frac{h(A_{\tau-\eta,\hat{\nu}},\hat{\nu})}{h(A_{\tau+\eta,\hat{\nu}},\hat{\nu})}\geq 1-e^{-\epsilon/2}\delta/3,$ (D.2)
then $(\operatorname{EM}_{\hat{\nu},\phi}(\cdot;\tau,\eta,\epsilon),\hat{\nu})$ is $(\epsilon/2,\delta)$-safe.
Proof.
This proof is inspired by the techniques of Brown et al., [6]. Let $Q=\operatorname{EM}_{\hat{\nu},\phi}(\cdot;\tau,\eta,\epsilon)$. For some $\tilde{\nu}\in\mathcal{M}(\hat{\nu},1)$, let $P=\operatorname{EM}_{\tilde{\nu},\phi}(\cdot;\tau,\eta,\epsilon)$. Take $C=A_{\tau,\tilde{\nu}}\cap A_{\tau,\hat{\nu}}$ and $E\in\mathscr{B}(\mathbb{R}^{d})$. Then,
$Q(E)$ $=Q(E\cap C)+Q(E\cap C^{c})$
$=\frac{Q(E\cap C)}{P(E\cap C)}P(E\cap C)+Q(E\cap A_{\tau,\tilde{\nu}}^{c})+Q(E\cap A_{\tau,\hat{\nu}}^{c})-Q(E\cap A_{\tau,\hat{\nu}}^{c}\cap A_{\tau,\tilde{\nu}}^{c})$
$\leq\frac{Q(E\cap C)}{P(E\cap C)}P(E\cap C)+Q(A_{\tau,\tilde{\nu}}^{c})$
$\leq\frac{Q(E\cap C)}{P(E\cap C)}P(E)+Q(A_{\tau,\tilde{\nu}}^{c}).$
Analogously, we have that
$P(E)$ $\leq\frac{P(E\cap C)}{Q(E\cap C)}Q(E\cap C)+P(A_{\tau,\hat{\nu}}^{c})\leq\frac{P(E\cap C)}{Q(E\cap C)}Q(E)+P(A_{\tau,\hat{\nu}}^{c}).$
It remains to show the following bounds:
$Q(A_{\tau,\tilde{\nu}}^{c}),P(A_{\tau,\hat{\nu}}^{c})$ $\leq\delta/3,\qquad\text{ and }\qquad\frac{Q(E\cap C)}{P(E\cap C)},\frac{P(E\cap C)}{Q(E\cap C)}\leq e^{\epsilon/2}+2\delta/3.$
First, (D.1) implies that for $\tilde{\nu}\in\mathcal{M}(\hat{\nu},1)$, it holds that both
$A_{\tau-\eta,\tilde{\nu}}\subset A_{\tau,\hat{\nu}}\subset A_{\tau+\eta,\tilde{\nu}}\qquad\text{and}\qquad A_{\tau-\eta,\hat{\nu}}\subset A_{\tau,\tilde{\nu}}\subset A_{\tau+\eta,\hat{\nu}}.$ (D.3)
In addition, noting that
[…]
then (D.1) implies that for any $A\subset A_{\tau,\tilde{\nu}}\cup A_{\tau,\hat{\nu}}$,
$e^{-\epsilon/4}h(A,\hat{\nu})\leq h(A,\tilde{\nu})\leq e^{\epsilon/4}h(A,\hat{\nu}).$ (D.4)
Applying (D.3) and (D.4) together with (D.2) yields
$\frac{Q(E\cap C)}{P(E\cap C)}$ $\leq e^{\epsilon/4}\frac{h(A_{\tau,\tilde{\nu}},\tilde{\nu})}{h(A_{\tau,\hat{\nu}},\hat{\nu})}\leq e^{\epsilon/2}\frac{h(A_{\tau,\tilde{\nu}},\hat{\nu})}{h(A_{\tau,\hat{\nu}},\hat{\nu})}\leq e^{\epsilon/2}\frac{h(A_{\tau+\eta,\hat{\nu}},\hat{\nu})}{h(A_{\tau,\hat{\nu}},\hat{\nu})}\leq e^{\epsilon/2}(1-e^{-\epsilon/2}\delta/3)^{-1}\leq e^{\epsilon/2}+2\delta/3,$
noting that for $x\leq 1/2$, $(1-x)^{-1}\leq 1+1/x$. By an analogous argument,
$\frac{P(E\cap C)}{Q(E\cap C)}$ $\leq e^{\epsilon/2}+2\delta/3.$
Next, applying (D.3) and (D.4) together with (D.2) yields
$Q(A_{\tau,\tilde{\nu}}^{c})$ $=1-\frac{h(A_{\tau,\tilde{\nu}}\cap A_{\tau,\hat{\nu}},\hat{\nu})}{h(A_{\tau,\hat{\nu}},\hat{\nu})}\leq 1-\frac{h(A_{\tau-\eta,\hat{\nu}},\hat{\nu})}{h(A_{\tau,\hat{\nu}},\hat{\nu})}\leq e^{-\epsilon/2}\delta/3\leq\delta/3.$
Similarly, we have that
$P(A_{\tau,\hat{\nu}}^{c})=\frac{h(A_{\tau,\tilde{\nu}}\cap A_{\tau,\hat{\nu}}^{c},\tilde{\nu})}{h(A_{\tau,\tilde{\nu}},\tilde{\nu})}$ $\leq\frac{e^{\epsilon/4}h(A_{\tau,\hat{\nu}}^{c}\cap A_{\tau,\tilde{\nu}},\hat{\nu})}{e^{-\epsilon/4}h(A_{\tau,\tilde{\nu}},\hat{\nu})}$
$\leq e^{\epsilon/2}\left(1-\frac{h(A_{\tau,\hat{\nu}}\cap A_{\tau,\tilde{\nu}},\hat{\nu})}{h(A_{\tau,\tilde{\nu}},\hat{\nu})}\right)$
$\leq e^{\epsilon/2}\left(1-\frac{h(A_{\tau-\eta,\hat{\nu}},\hat{\nu})}{h(A_{\tau+\eta,\hat{\nu}},\hat{\nu})}\right)$
$\leq e^{\epsilon/2}\left(e^{-\epsilon/2}\delta/3\right)=\delta/3.\qed$
Lemma D.2.
If, for $\hat{\nu}\in\widehat{\mathcal{M}}_{1}(n,d)$, $k\in[n]$, $\epsilon>0$, $20<\delta\leq e^{-\epsilon/2}/$ 2 and $\operatorname{EM}_{\hat{\nu},\phi}=\operatorname{EM}_{\hat{\nu},\phi}(\cdot;\tau,\eta,\epsilon)$ it holds that
$\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|\phi(x,\hat{\nu})-\phi(x,\tilde{\nu})|\leq\eta/2$ (D.5)
and
$\inf_{y>0}e^{-\epsilon y/4\eta}\frac{\operatorname{vol}(A_{\tau+2\eta,\hat{\nu}}\cap A_{\tau-2\eta,\hat{\nu}}^{c})}{\operatorname{vol}(A_{\tau-y-2\eta,\hat{\nu}})}\leq e^{-\epsilon/4}\delta/3,$ (D.6)
then $\operatorname{SM}(\hat{\nu},\operatorname{EM}_{\hat{\nu},\phi},\epsilon/2,\delta)\geq k-1$.
Proof.
To say that $\operatorname{SM}(\hat{\nu},\operatorname{EM}_{\hat{\nu},\phi},\epsilon/2,\delta)\geq k-1$ is equivalent to saying that, for every $\tilde{\nu}\in\mathcal{M}(\hat{\nu},k-1)$, $(\operatorname{EM}_{\tilde{\nu},\phi},\tilde{\nu})$ is $(\epsilon/2,\delta)$-safe. Using Lemma D.1 on $\tilde{\nu}$, we need only show that (D.1) and (D.2) hold. That is
$\sup_{\begin{subarray}{c}\nu^{\prime}\in\widetilde{\mathcal{M}}_{1}(\tilde{\nu},1)\\ x\in A_{\tau,\tilde{\nu}}\cup A_{\tau,\nu^{\prime}}\end{subarray}}|\phi(x,\tilde{\nu})-\phi(x,\nu^{\prime})|\leq\eta,\qquad\text{and}\qquad\frac{h(A_{\tau-\eta,\tilde{\nu}},\tilde{\nu})}{h(A_{\tau+\eta,\tilde{\nu}},\tilde{\nu})}\geq 1-e^{-\epsilon/2}\delta/3.$
First, the assumption (D.5), together with the triangle inequality, implies that, for every $v^{\prime}\in\widetilde{\mathcal{M}}_{1}(\tilde{\nu},1)$ and $x\in A_{\tau+\eta,\tilde{\nu}}\cup A_{\tau+\eta,\nu^{\prime}}$,
$|\phi(x,\nu^{\prime})-\phi(x,\tilde{\nu})|=|\phi(x,\nu^{\prime})-\phi(x,\hat{\nu})+\phi(x,\hat{\nu})-\phi(x,\tilde{\nu})|\\ \leq 2\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|\phi(x,\hat{\nu})-\phi(x,\tilde{\nu})|\leq\eta.$
This, taken together with the fact that $A_{\tau,\nu}\subseteq A_{\tau+\eta,\nu}$ for every $\nu$, implies that (D.1) holds.
Next, note that
$\frac{h(A_{\tau-\eta,\tilde{\nu}},\tilde{\nu})}{h(A_{\tau+\eta,\tilde{\nu}},\tilde{\nu})}=\frac{h(A_{\tau-\eta,\tilde{\nu}}\cap A_{\tau+\eta,\tilde{\nu}},\tilde{\nu})}{h(A_{\tau+\eta,\tilde{\nu}},\tilde{\nu})}=\operatorname{EM}_{\tilde{\nu},\phi}(A_{\tau-\eta,\tilde{\nu}};\tau+\eta,\eta,\epsilon).$
As a result, (D.2) is equivalent to
$\operatorname{EM}_{\tilde{\nu},\phi}(A_{\tau-\eta,\tilde{\nu}};\tau+\eta,\eta,\epsilon)\geq 1-e^{-\epsilon/2}\delta/3\implies e^{-\epsilon/2}\delta/3\geq 1-\operatorname{EM}_{\tilde{\nu},\phi}(A_{\tau-\eta,\tilde{\nu}};\tau+\eta,\eta,\epsilon).$
Thus, we turn to proving the condition
$31-\operatorname{EM}_{\tilde{\nu},\phi}(A_{\tau-\eta,\tilde{\nu}};\tau+\eta,\eta,\epsilon)=\frac{h(A_{\tau-\eta,\tilde{\nu}}^{c}\cap A_{\tau+\eta,\tilde{\nu}},\tilde{\nu})}{h(A_{\tau+\eta,\tilde{\nu}},\tilde{\nu})}\leq e^{-\epsilon/2}\delta/$ 3 . (D.7)
To this end, note that for any $A\subset A_{\tau+\eta,\tilde{\nu}}\cup A_{\tau+\eta,\hat{\nu}}$,
$e^{-\epsilon/4}h(A,\nu^{\prime})\leq h(A,\tilde{\nu})\leq e^{\epsilon/4}h(A,\nu^{\prime}).$
Moreover, (D.5) implies that $A_{\tau,\tilde{\nu}}\subset A_{\tau+\eta,\nu^{\prime}}\subset A_{\tau+2\eta,\tilde{\nu}}$ and $A_{\tau,\nu^{\prime}}\subset A_{\tau+\eta,\tilde{\nu}}\subset A_{\tau+2\eta,\nu^{\prime}}$. (D.3) Applying these two facts yields that for any $y>0$, it holds that
$h(A_{\tau+\eta,\tilde{\nu}},\tilde{\nu})\geq h(A_{\tau-y-2\eta,\tilde{\nu}},\tilde{\nu})\geq e^{-\epsilon/4}h(A_{\tau-y-2\eta,\nu^{\prime}},\nu^{\prime})\geq e^{-\epsilon(\tau-y-\eta)/(4\eta)}\operatorname{vol}(A_{\tau-y-2\eta,\nu^{\prime}}).$ (D.8)
Next, we bound the numerator in the left-hand side of (D.7) using the same two facts,
$h(A_{\tau+\eta,\tilde{\nu}}\cap A_{\tau-\eta,\tilde{\nu}}^{c},\tilde{\nu})\leq e^{\epsilon/4}h(A_{\tau+2\eta,\hat{\nu}}\cap A_{\tau-2\eta,\hat{\nu}}^{c},\hat{\nu})\leq e^{-\epsilon(\tau+\eta)/(4\eta)}\operatorname{vol}(A_{\tau+2\eta,\hat{\nu}}\cap A_{\tau-2\eta,\hat{\nu}}^{c}).$ (D.9)
Lastly, combining (D.8) and (D.9) with (D.6) yields
$\frac{h(A_{\tau-\eta,\tilde{\nu}}^{c}\cap A_{\tau+\eta,\tilde{\nu}},\tilde{\nu})}{h(A_{\tau+\eta,\tilde{\nu}},\tilde{\nu})}$ $\leq e^{-\epsilon(y+\eta)/4\eta}\frac{\operatorname{vol}(A_{\tau+2\eta,\hat{\nu}}\cap A_{\tau-2\eta,\hat{\nu}}^{c})}{\operatorname{vol}(A_{\tau-y-2\eta,\hat{\nu}})}\leq e^{-\epsilon/2}\delta/3,$
which completes the proof.
Proof of Lemma 5.3.
Using the fact that $W$ is a standard Laplace random variable, for all $k\in\{0\}\cup[n-1]$, it holds that
$1-\mathbb{E}\lambda(\hat{\nu},\epsilon,\delta)$ $=\Pr\left(\operatorname{SM}(\hat{\nu},Q)+\frac{2}{\epsilon}W<\frac{2\log(1/2\delta)}{\epsilon}\right)$
$\leq\Pr\left(\operatorname{SM}(\hat{\nu},Q)<k\right)+\Pr\left(W<\log(1/2\delta)-\epsilon k/2\right).$
It is immediate from the fact that the above bound holds for all $k\in\{0\}\cup[n-1]$ that
$1-\mathbb{E}\lambda(\hat{\nu},\epsilon,\delta)$ $\leq\inf_{k\in\{0\}\cup[n-1]}\left[\Pr\left(\operatorname{SM}(\hat{\nu},Q)<k\right)+\Pr\left(W<\log(1/2\delta)-\epsilon k/2\right)\right].$
Applying Lemma D.2 yields that, for all $k\in\{0\}\cup[n-1]$, it holds that
$\Pr\left(\operatorname{SM}(\hat{\nu},Q)<k\right)\leq\Pr\Big{(}\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k+1)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|\phi(x,\hat{\nu})-\phi(x,\tilde{\nu})|>\eta/2\Big{)}\\ +\Pr\left(\inf_{y>0}e^{-\epsilon y/4\eta}\frac{\operatorname{vol}(A_{\tau+2\eta,\hat{\nu}}\cap A_{\tau-2\eta,\hat{\nu}}^{c})}{\operatorname{vol}(A_{\tau-y-2\eta,\hat{\nu}})}>e^{-\epsilon/4}\delta/3\right).$
Combining the two bounds gives the result. ∎
Appendix E Results concerning Condition 2
Here, we show that the median and median absolute deviation satisfy Condition 2 under general conditions.
Lemma E.1.
Let $\mu$ be the median. If Condition 4 holds for $(1/2,\nu,0)$ then for all $t\geq 0$ and $n,d\geq 1$
$\Pr\left(S_{n}(\mu)\geq t\right)\lesssim\exp\left(-2nM^{2}t^{2}+d\log(cn/d)\right).$
Proof.
The proof follows from Lemma C.1. ∎
Lemma E.2.
Let $\sigma$ be the median absolute deviation. If for some $q^{\prime}$ that satisfies
$0<q^{\prime}<\inf_{u\in\mathbb{S}^{d-1}}\operatorname*{{\rm argmin}}_{0<p\leq 1/2}|\xi_{1/2,u}-\xi_{p,u}-(\xi_{p+1/2,u}-\xi_{1/2,u})|,$
Condition 4 holds for $(q^{\prime},\nu,0)$ with $r\geq 1/3M$, and $\sup_{\begin{subarray}{c}x\in\mathcal{I};u\in\mathbb{S}^{d-1}\end{subarray}}f_{\nu_{u}}(x)<L<\infty$, then for all $t\geq 0$ and $n,d\geq 1$
$\Pr\left(S_{n}(\sigma)\geq t\right)\lesssim\exp\left(-2nM^{2}t^{2}/16^{2}\vee-2n[(1/2-q^{\prime})M/L\wedge 1/3]^{2}+d\log(cn/d)\right).$
Proof.
Let
$\mathcal{C}_{u}(p)=\xi_{1/2,u}-\xi_{p,u}-(\xi_{p+1/2,u}-\xi_{1/2,u}),$
and $q_{u}=\operatorname*{{\rm argmin}}_{0<p\leq 1/2}|\mathcal{C}_{u}(p)|$. Now, note that under absolute continuity of $\nu_{u}$, it holds that $\sigma(\nu_{u})=|\xi_{1/2,u}-\xi_{q_{u},u}|$. This is because the median absolute deviation of a measure $P$ is half the length of the symmetric interval about the median, which contains 1/2 of the probability mass of $P$ [21]. Now, analogously, let
$\hat{\mathcal{C}}_{u}(p)=\hat{\xi}_{1/2,u}-\hat{\xi}_{p,u}-(\hat{\xi}_{p+1/2,u}-\hat{\xi}_{1/2,u})\qquad\text{and}\qquad\tilde{\mathcal{C}}_{u}(p)=\hat{\xi}_{1/2,u}-\hat{\xi}_{p-1/2,u}-(\hat{\xi}_{p,u}-\hat{\xi}_{1/2,u}),$
with
$\hat{q}_{u}=\operatorname*{{\rm argmin}}_{0\leq p\leq 1/2}|\hat{\mathcal{C}}_{u}(p)|\qquad\text{and}\qquad\tilde{q}_{u}=\operatorname*{{\rm argmin}}_{1/2\leq p\leq 1}|\tilde{\mathcal{C}}_{u}(p)|.$
It follows by definition of the median absolute deviation that for any $u\in\mathbb{S}^{d-1}$,
$|\hat{\xi}_{1/2,u}-\hat{\xi}_{\hat{q}_{u},u}|\wedge|\hat{\xi}_{1/2,u}-\hat{\xi}_{\tilde{q}_{u},u}|\leq\sigma(\hat{\nu}_{u})\leq|\hat{\xi}_{1/2,u}-\hat{\xi}_{\hat{q}_{u},u}|\vee|\hat{\xi}_{1/2,u}-\hat{\xi}_{\tilde{q}_{u},u}|.$
This fact directly implies that for any $u\in\mathbb{S}^{d-1}$, either
$|\sigma(\hat{\nu}_{u})-\sigma(\nu_{u})|\leq|\xi_{1/2,u}-\xi_{q_{u},u}-(\hat{\xi}_{1/2,u}-\hat{\xi}_{\hat{q}_{u},u})|\leq|\xi_{1/2,u}-\hat{\xi}_{1/2,u}|+|\xi_{\hat{q}_{u},u}-\xi_{q_{u},u}|+|\xi_{\hat{q}_{u},u}-\hat{\xi}_{\hat{q}_{u},u}|,$ (E.1)
holds, or
$|\sigma(\hat{\nu}_{u})-\sigma(\nu_{u})|\leq|\xi_{1/2,u}-\hat{\xi}_{1/2,u}|+|\xi_{\tilde{q}_{u},u}-\xi_{q_{u}+1/2,u}|+|\xi_{\tilde{q}_{u},u}-\hat{\xi}_{\tilde{q}_{u},u}|,$ (E.2)
holds. Next, we bound $\hat{q}_{u}$ below and $\tilde{q}_{u}$ above for all $u\in\mathbb{S}^{d-1}$. Consider $w\in\mathbb{R}$ such that $q^{\prime}<w<\inf_{u\in\mathbb{S}^{d-1}}q_{u}$. Now, it follows from Condition 4 and an application of the mean value theorem that
$\xi_{w,u}-\xi_{q^{\prime},u}\leq(w-q^{\prime})/\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ x\in(\xi_{q^{\prime},u},\xi_{w,u})\end{subarray}}f_{u}(x)\leq M^{-1}.$
Now, using Condition 4 and the assumption that $r>1/3M$, together with an application of Lemma C.1 yields
$\Pr\Big{(}\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ x\in\{w,1/2\}\end{subarray}}|\hat{\xi}_{x,u}-\xi_{x,u}|>1/3M\Big{)}\lesssim\left(\frac{cn}{d}\right)^{d}e^{-2n/9}.$ (E.3)
Next, we show that the probability that $(\hat{\xi}_{w,u},\hat{\xi}_{1/2,u}+(\hat{\xi}_{1/2,u}-\hat{\xi}_{w,u}))=(\hat{\xi}_{w,u},2\hat{\xi}_{1/2,u}-\hat{\xi}_{w,u})$ contains less than 1/2 of the projected observations for at least one $u\in\mathbb{S}^{d-1}$ is small. This immediately implies that $\hat{q}_{u}\geq w>q^{\prime}$ for any $u\in\mathbb{S}^{d-1}$. To this end, (E.3) implies that with high probability,
$(\hat{\xi}_{w,u},2\hat{\xi}_{1/2,u}-\hat{\xi}_{w,u})\supset(\xi_{w}+1/3M,2\xi_{1/2,u}-\xi_{w}-1/M)\supset(\xi_{q^{\prime}},2\xi_{1/2}-\xi_{q^{\prime}}).$
Now, we bound the probability that $(\xi_{q^{\prime},u},2\xi_{1/2,u}-\xi_{q^{\prime},u})$ contains less than 1/2 of the projected observations. This is equivalent to bounding:
$\Pr\Big{(}\bigcup_{u\in\mathbb{S}^{d-1}}\{\hat{\xi}_{1/2,u}<\xi_{q^{\prime},u}\}\Big{)}\qquad\text{and}\qquad\Pr\Big{(}\bigcup_{u\in\mathbb{S}^{d-1}}\{\hat{\xi}_{1/2,u}>2\xi_{1/2,u}-\xi_{q^{\prime},u}\}\Big{)}.$ (E.4)
Starting with the left-hand term, utilizing the mean value theorem, together with the upper bound on the projected density and an application of Lemma C.1, yields
$\Pr\Big{(}\bigcup_{u\in\mathbb{S}^{d-1}}\{\hat{\xi}_{1/2,u}<\xi_{q^{\prime},u}\}\Big{)}$ $\leq\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}|\hat{\xi}_{1/2,u}-\xi_{1/2,u}|>\inf_{u\in\mathbb{S}^{d-1}}[\xi_{1/2,u}-\xi_{q^{\prime},u}]\right)$
$\leq\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}|\hat{\xi}_{1/2,u}-\xi_{1/2,u}|>(1/2-q^{\prime})/L\right)$
$\leq\left(\frac{cn}{d}\right)^{d}e^{-2nM^{2}(1/2-q^{\prime})^{2}/L^{2}}.$
Similarly, the right-hand term in (E.4) satisfies
$\Pr\Big{(}\bigcup_{u\in\mathbb{S}^{d-1}}\{\hat{\xi}_{1/2,u}>\xi_{1/2,u}+\xi_{q^{\prime},u}\}\Big{)}$ $\leq\left(\frac{cn}{d}\right)^{d}e^{-2nM^{2}(1/2-q^{\prime})^{2}/L^{2}}.$
Thus,
$\Pr\left(\inf_{u\in\mathbb{S}^{d-1}}\hat{q}_{u}<q^{\prime}\right)\lesssim\left(\frac{cn}{d}\right)^{d}e^{-2nM^{2}[(1/2-q^{\prime})/L\wedge M/3]^{2}}.$
An analogous argument gives that
$\Pr\Big{(}\sup_{u\in\mathbb{S}^{d-1}}\tilde{q}_{u}>1-q^{\prime}\Big{)}\lesssim\left(\frac{cn}{d}\right)^{d}e^{-2nM^{2}[(1/2-q^{\prime})/L\wedge M/3]^{2}}.$
Let $E_{11}=\{q^{\prime}<\hat{q}_{u}<\tilde{q}_{u}\leq 1-q^{\prime}\}$ and note that we have just shown that
$\Pr\left(E_{11}^{c}\right)\lesssim\left(\frac{cn}{d}\right)^{d}e^{-2nM^{2}[(1/2-q^{\prime})/L\wedge M/3]^{2}}.$ (E.5)
The next step is to show that $\{\sup_{u\in\mathbb{S}^{d-1}}|\xi_{\hat{q},u}-\xi_{q,u}|\geq t|E_{11}\}\subset\{\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ q^{\prime}<p<1-q^{\prime}\end{subarray}}|\xi_{p,u}-\hat{\xi}_{p,u}|\gtrsim t|E_{11}\}$. To this end, observe that
$\{\sup_{u\in\mathbb{S}^{d-1}}|\xi_{\hat{q},u}-\xi_{q,u}|\geq t|E_{11}\}$ $=\bigcup_{u\in\mathbb{S}^{d-1}}\{\inf_{\begin{subarray}{c}q^{\prime}<p<1/2\\ |\xi_{p,u}-\xi_{q,u}|>t\end{subarray}}|\hat{\mathcal{C}}_{u}(p)|-|\hat{\mathcal{C}}_{u}(q)|\leq 0\}$
$=\bigcup_{u\in\mathbb{S}^{d-1}}\{|\hat{\mathcal{C}}_{u}(q)|-\inf_{\begin{subarray}{c}q^{\prime}<p<1/2\\ |\xi_{p,u}-\xi_{q,u}|>t\end{subarray}}|\hat{\mathcal{C}}_{u}(p)|\geq 0\}$
$\subset\{8\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ q^{\prime}<p<1-q^{\prime}\end{subarray}}|\xi_{p,u}-\hat{\xi}_{p,u}|\geq\inf_{p<1/2,\ |\xi_{p,u}-\xi_{q,u}|>t}|\mathcal{C}_{u}(p)|\}.$
It remains to show that for any $u\in\mathbb{S}^{d-1}$, $\inf_{0<p<1/2,\ |\xi_{p,u}-\xi_{q,u}|>t}|\mathcal{C}_{u}(p)|\gtrsim t$. This is equivalent to
$\inf_{\begin{subarray}{c}0<p<1/2\\ |\xi_{p,u}-\xi_{q,u}|>t\end{subarray}}|\xi_{q,u}-\xi_{p,u}+\xi_{q+1/2,u}-\xi_{p+1/2,u}|\gtrsim t.$
Now,
$\inf_{\begin{subarray}{c}0<p<1/2\\ |\xi_{p,u}-\xi_{q,u}|>t\end{subarray}}|\xi_{q,u}-\xi_{p,u}+\xi_{q+1/2,u}-\xi_{p+1/2,u}|=\inf_{\begin{subarray}{c}0<p<1/2\\ |\xi_{p,u}-\xi_{q,u}|>t\end{subarray}}[|\xi_{q,u}-\xi_{p,u}|+|\xi_{q+1/2,u}-\xi_{p+1/2,u}|]\geq t.$
The first equality comes from the fact that $|\xi_{p,u}-\xi_{q,u}+\xi_{q+1/2,u}-\xi_{p+1/2,u}|$ is of the form of $|a+b|$ where $sign(a)=sign(b)$. The second inequality comes from the fact that we are minimizing over $|\xi_{p,u}-\xi_{q,u}|>t$ and the fact that the second term is non-negative. Thus,
$\{|\xi_{\hat{q},u}-\xi_{q,u}|\geq t|E_{11}\}\subset\{\sup_{q^{\prime}<p<1-q^{\prime}}|\xi_{p,u}-\hat{\xi}_{p,u}|\geq t/8\}.$ (E.6)
By an analogous argument,
$\{|\xi_{q,u}-\xi_{\tilde{q},u}|\geq t|E_{11}\}\subset\{\sup_{q^{\prime}<p<1-q^{\prime}}|\xi_{p,u}-\hat{\xi}_{p,u}|\geq t/8\}.$ (E.7)
Now, combining equations (E.1), (E.2), (E.5), (E.6) and (E.7),
$\Pr\left(S_{n}(\sigma)\geq t\right)$ $\lesssim\Pr\left(\sup_{q^{\prime}<p<1-q^{\prime}}|\xi_{p,u}-\hat{\xi}_{p,u}|\geq t/16\right)+\left(\frac{cn}{d}\right)^{d}e^{-2n[(1/2-q^{\prime})M/L\wedge 1/3]^{2}}$
$\lesssim\left(\frac{cn}{d}\right)^{d}e^{-2nM^{2}t^{2}/16^{2}}+\left(\frac{cn}{d}\right)^{d}e^{-2n[(1/2-q^{\prime})M/L\wedge 1/3]^{2}}.\qed$
o1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">